### 02 Apr 2026 v1

# Hölder regularity of Siciak-Zaharjuta extremal functions on compact Hermitian manifolds

Abstract.
For a compact subset in a compact Hermitian manifold, we prove that the Hölder continuity of the extremal function at a given point in the set is a local property and that the Hölder continuity of a weighted extremal function follows from the Hölder continuities of the extremal function and the weight function with a uniform density in capacity. The second result can be seen as a continuation of a result of Lu, Phung and Tô [LPT21]. Moreover, for a compact subset in a compact Hermitian manifold, we prove that the Hölder continuity of the extremal function with the uniform density in capacity is equivalent to the local Hölder continuity property. These results are generalizations of the results of Nguyen [Ng24] on compact Kähler manifolds. We also show that the $\mu$-Hölder continuity property of a convex compact subset in $\mathbb{C}^{n}$ implies the local $\mu$-Hölder continuity property of order $1$.
Key words and phrases:
extremal function, Hermitian manifold, Hölder continuous, uniform density in capacity, HCP, local Hölder continuity property
1. Introduction
Background. Complex Monge-Ampère equations are essential for the classification of complex surfaces and in solving geometric flow problems [PSS12]. Solving complex Monge-Ampère equations on compact Hermitian manifolds is important in pluripotential theory because it generalizes Kähler geometry to broader non-Kähler settings and provides a framework for analyzing degenerate equations [GZ17]. Stability and the Hölder regularity of the solutions to the equations on compact Hermitian manifolds are obtained in [KN18].
The Hölder continuity of the solution to the complex Monge-Ampère equations is the best global regularity when the density is in $L^{p}$ with $p>1$. (Let $\|\cdot\|$ be the standard Euclidean norm on $\mathbb{C}^{n}$. For each $p\in(1,2)$ and $\beta\in(2-2/p,1)$, $u(z):=\|z\|^{\beta}$ has $L^{p}$-density, but $u$ is not Lipschitz continuous at $\mathbf{0}$.) Kołodziej [Ko03] first proved the Hölder continuity with an unspecified Hölder exponent. After that, Demailly, Dinew, Guedj, Pham, Kołodziej and Zeriahi in [DDGHKZ] showed it with an explicit exponent on compact Kähler manifolds. After that, Kołodziej and Nguyen in [KN19] showed the Hölder continuity on compact Hermitian manifolds with strictly positive density. Also, Kołodziej and Nguyen in [KN18] showed it on compact Hermitian manifolds with non-negative density.
In fact, Kołodziej and Nguyen showed more than that. They [KN18] proved that on a compact Hermitian manifold $(X,\omega)$, complex Monge-Ampère equations with finite Borel measures, dominated by complex Monge-Ampère measures of $\omega$-plurisubharmonic Hölder continuous functions (in other words, admitting global Hölder continuous subsolutions), have Hölder continuous solutions. Lu, Phung and Tô in [LPT21] showed a nearly optimal Hölder exponent of the solutions to the equations in compact Hermitian manifolds, as in compact Kähler manifolds.
The Siciak-Zaharjuta extremal function can serve as such a subsolution when it is Hölder continuous. Lu, Phung and Tô in [LPT21] also showed that for a Hölder continuous weight function on a compact Hermitian manifold, its $\omega$-plurisubharmonic envelope becomes an $\omega$-plurisubharmonic Hölder continuous function. We continue the result of [LPT21] for general compact subsets of compact Hermitian manifolds.
Recall that the Lelong class $\mathcal{L}$ is defined by
$\mathcal{L}:=\{u\in PSH(\mathbb{C}^{n}):\text{for some constant }c_{u},\;u(z)\leq\max\{0,\log\|z\|\}+c_{u},\;z\in\mathbb{C}^{n}\}.$
For a subset $E$ in $\mathbb{C}^{n}$, the classical Siciak-Zaharjuta extremal function of $E$ introduced in [Si62, Si81, Za76] is defined on $\mathbb{C}^{n}$ by
$L_{E}(z):=\sup\{u(z):u\in\mathcal{L},\,u|_{E}\leq 0\},\quad z\in\mathbb{C}^{n}.$
The regularity of the extremal function of a set gives geometric information about the set itself. For example, $E$ is pluripolar if and only if $L_{E}^{*}\equiv\infty$, where denotes the upper semicontinuous regularization, and $E$ is non-pluripolar if and only if $L_{E}^{*}\in\mathcal{L}$. For a real-valued function $\phi$ on $E$, the weighted extremal function $L_{E,\phi}$ is defined by
$L_{E,\phi}(z):=\sup\{u(z):u\in\mathcal{L},u|_{E}\leq\phi\},\quad z\in\mathbb{C}^{n}.$
Results. Throughout this article, $(X,\omega)$ is a compact Hermitian manifold of complex dimension $n$. A function $\phi:X\to\mathbb{R}\cup\{-\infty\}$ is called quasi-plurisubharmonic (quasi-psh for short) if $\phi\not\equiv$-$\infty$ and $\phi$ is locally written as the sum of a smooth function and a psh function. For the real differential operators $d:=\partial+\bar{\partial}$ and $d^{c}:=i(\bar{\partial}-\partial)$, the family of $\omega$-plurisubharmonic ($\omega$-psh for short) functions on $X$ is defined by
$PSH(X,\omega):=\{v:X\to\mathbb{R}\cup\{-\infty\}:v\text{ is quasi-psh},\,\omega+dd^{c}v\geq 0\}.$
If $\phi:X\to\mathbb{R}\cup\{-\infty\}$ is quasi-psh, then there exists a constant $c>0$ satisfying $dd^{c}\phi+c\omega\geq 0$, since $X$ is compact and $PSH(X,c_{1}\omega)\subset PSH(X,c_{2}\omega)$ for $0<c_{1}<c_{2}<\infty$.
The extremal function $V_{E}=V_{\omega;E}$ of a subset $E$ in $X$ is defined by
(1.1) $V_{E}(z)=V_{\omega;E}(z):=\sup\{v(z):v\in PSH(X,\omega):v|_{E}\leq 0\}.$
For a weight function $\phi:E\to\mathbb{R}$, the weighted extremal function of $(E,\phi)$ is
(1.2) $V_{E,\phi}(z)=V_{\omega;E,\phi}(z):=\sup\{v(z):v\in PSH(X,\omega),\;v|_{E}\leq\phi\}.$
When the domain of a real-valued function $\psi$ contains $E$, we write $V_{E,\psi|_{E}}$ as $V_{E,\psi}$ for convenience. There is a natural one-to-one correspondence between the Lelong class $\mathcal{L}$ and $PSH(\mathbb{CP}^{n},\omega_{FS})$ ([GZ17, Chapter 8]). When $\phi$ is a continuous function on a compact subset of $X$, we can extend $\phi$ to a continuous function on $X$ with the same supremum norm by Tietze’s extension theorem. An extension of a Hölder continuous function from a compact subset to $X$ is also possible with the same Hölder exponent and the same Hölder constant by Mcshane’s extension [Mc34].
Our main theorems are about the Hölder continuity of extremal functions for non-pluripolar compact sets in $X$. Lu, Phung and Tô in [LPT21] proved that for a given Hölder continuous real-valued function $\phi$ on $X$ with a Hölder exponent $\alpha\in(0,1)$, the envelope $V_{X,\phi}^{*}$ is also Hölder continuous with the same exponent $\alpha$. $V_{X,\phi}$ is continuous by [Ahn26a] since $V_{X}\equiv 0$ and $\phi$ are continuous, thus $V_{X,\phi}^{*}=V_{X,\phi}$. However, the Hölder exponent or even Hölder continuity of the weighted extremal functions for a general non-pluripolar compact subset is not always preserved. We need an additional condition, called the uniform density in capacity [Ng24].
Let $\Omega$ be a smoothly bounded strictly pseudoconvex domain in $\mathbb{C}^{n}$. Bedford and Taylor in [BT82] first studied the relative capacity of a Borel subset $E$ in $\Omega$, defined by
$Cap(E,\Omega):=\sup\{\int_{E}(dd^{c}u)^{n}:u\in PSH(\Omega),\,0\leq u\leq 1\}.$
We denote the Euclidean open ball of center $a\in\mathbb{C}^{n}$ and radius $r\in(0,\infty]$ by $\mathbb{B}(a,r)$ and its closure by $\bar{\mathbb{B}}(a,r)$. Let $F$ be a compact subset of $\mathbb{C}^{n}$. We say that $F$ has a uniform density in capacity if there exist constants $q>0$, $\varkappa>0$, $\delta>0$ and a Euclidean bounded open ball $O$ in $\mathbb{C}^{n}$ such that $\bar{\mathbb{B}}(F,\delta)$ is contained in $O$ and for each $b\in F$,
(1.3) $\inf_{0<r\leq\delta}\frac{Cap(F\cap\bar{\mathbb{B}}(b,r),O)}{r^{q}}\geq\varkappa.$
We say that $F$ has a uniform density in capacity at $b\in F$ if (1.3) holds. Recall that the global Bedford-Taylor capacity of a Borel subset $E$ in $X$ is defined as
(1.4) $Cap_{\omega}(E):=\sup\{\int_{E}(\omega+dd^{c}v)^{n}:v\in PSH(X,\omega),\,0\leq v\leq 1\}.$
By [DK12, Proposition 2.3], $Cap_{\omega}(X)<\infty$. The property of the uniform density in capacity of a compact subset in $X$ is defined in Section 3 using the global Bedford-Taylor capacity, similarly as that of a compact subset in $\mathbb{C}^{n}$.
Our main results are about the Hölder continuity of extremal functions on $X$.
Theorem 1.1.
Let $K$ be a compact subset of a compact Hermitian manifold $X$ and $a\in K$. Let $\bar{B}(a,r)$ be a closed holomorphic coordinate ball in $X$ with the center $a\in K$ and the finite radius $r>0$. Then the following items hold:
  • (i)
    Let $\mu\in(0,1]$. $V_{K}$ is $\mu$-Hölder continuous at $a$ if and only if $V_{K\cap\bar{B}(a,r)}$ is $\mu$-Hölder continuous at $a$.
  • (ii)
    If $K$ has a uniform density in capacity and $\phi:X\to\mathbb{R}$ and $V_{K}$ are Hölder continuous, then $V_{K,\phi}$ is Hölder continuous.
Theorem 1.2.
Let $K$ be a compact subset of a compact Hermitian manifold $X$ and $a\in K$. Then the following three items are equivalent.
  • (i)
    $V_{K}$ is Hölder continuous at $a$ and $K$ has a uniform density in capacity at $a$.
  • (ii)
    $K$ has local Hölder continuity property at $a$.
  • (iii)
    $K$ has weak local Hölder continuity property at $a$.
Also, the following three items are equivalent.
  • (iv)
    $V_{K}$ is Hölder continuous and $K$ has a uniform density in capacity.
  • (v)
    $K$ has local Hölder continuity property.
  • (vi)
    $K$ has weak local Hölder continuity property.
In Theorem 1.2, the equivalence of (iv), (v) and (vi) does not directly comes from that of (i), (ii) and (iii). The proof of Theorem 1.1-(i) gives Corollary 5.1, which is used for the proofs of Theorem 1.1-(ii) and Theorem 1.2. The local Hölder continuity property and the weak local Hölder continuity property in Theorem 1.2 are introduced in Section 3. Theorem 1.1-(ii) is a generalization of the above-mentioned result of [LPT21] to a general weighted extremal function, with the additional condition uniform density in capacity. However, unlike the envelopes, the Hölder exponent becomes smaller: if $V_{K}$ and $\phi$ are $\mu$-Hölder continuous for $\mu\in(0,1]$ and $K$ has a uniform density in capacity of order $q\in(0,\infty)$, then $V_{K,\phi}$ is $\mu^{2}/(\mu+q+2)$-Hölder continuous. N. C. Nguyen in [Ng24] proved Theorem 1.1 and some analogue of Theorem 1.2 for compact Kähler manifolds, and we generalize his results to compact Hermitian manifolds.
Let $0<\mu,\nu\leq 1$ and $0<q,q_{0}$ be constants. In Theorem 1.2, weak local $\mu$-HCP of order $q$ implies a uniform density in capacity of order $nq$ and $2\mu/(q+2)$-Hölder continuity. Conversely, $\nu$-Hölder continuity and a uniform density in capacity of order $q_{0}$ imply local $\nu$-HCP of order $2q_{0}+2$. Also, $\mu$-HCP of order $q$ implies weak local $\mu$-HCP of order $q$ by definitions. Conversely, in Remark 5.4, we give the proof that weak local $\mu$-HCP of order $q$ directly implies local $\mu$-HCP of order $(n+1)q$.
Remark 1.3.
There are several compact subsets in $\mathbb{C}^{n}$ with uniform densities in capacity. Since $\lim_{r\to 0^{+}}r^{\epsilon}\log{r}=0$ for each $\epsilon\in(0,\infty)$, closed Euclidean balls with finite radii have uniform densities in capacities by [Kl91, Example 5.1.1] and [GZ17, Example 4.37]. The image of a compact subset in $\mathbb{C}^{n}$ with a uniform density in capacity by an automorphism of $\mathbb{C}^{n}$ also has a uniform density in capacity. Compact subsets of $\mathbb{C}^{n}$ with local Hölder continuity property have uniform densities in capacity by Lemma 3.5, which is [Ng24, Lemma 3.3] proved by Nguyen. In [Ng24, Example 4.5], he gives several compact subsets in $\mathbb{C}^{n}$ with local Hölder continuity property, thus they have uniform densities in capacity.
For $\mathbb{K}=\mathbb{R}$ or $\mathbb{K}=\mathbb{C}$, the closure of a bounded Lipschitz domain in $\mathbb{K}^{n}$ and the closure of a bounded domain in $\mathbb{K}^{n}$ with the geometrical condition are shown to have the local Hölder continuity property as subsets of $\mathbb{C}^{n}$ in [Ng24, Example 4.5]. In particular, the real cube $0,1]^{n}$, the polydisk $\bar{\mathbb{B}}(\mathbf{0},1)^{n}$ and a convex compact subset of $\mathbb{K}^{n}$ with the nonempty interior in $\mathbb{K}^{n}$ has a uniform density in capacity in $\mathbb{C}^{n}$. In [Ng24, Example 4.5], the uniform interior sphere condition implies the geometrical condition since the union of uniform closed balls conjecture was proved by Nour, Stern and Takche [NST11]. [Ng24, Lemma 4.7-(a)] directly showed that the closure of a bounded domain in $\mathbb{C}^{n}$ with a uniform interior sphere condition implies the local Hölder continuity property with improvement in both the exponent and the order compared to the geometric condition. This direct proof is also valid since the uniform interior $r$-sphere condition implies the union of uniform closed $r/2$-balls by [NST11].
Remark 1.4.
The main difficulty of Theorem 1.2 compared to the case for compact Kähler manifolds is the inequality (2.5), which compares the supremum of the extremal function and the global Bedford-Taylor capacity on $X$. Whether a Hermitian metric $\omega$ on $X$ is uniformly non-collapsing remains an open problem. Although Guedj and Lu [GL22, Proposition 3.4] proved that
$0<\inf\{\int_{X}(\omega+dd^{c}v)^{n}:v\in PSH(X,\omega),\;0\leq v\leq M\}$
for each $M\in(0,\infty)$, but an explicit uniform lower bound for the infimum in terms of $X,\omega$ and $M$ is not known. To avoid this difficulty, we compare the supremum of the extremal function and the relative capacity in $\mathbb{C}^{n}$ to prove Theorem 1.2.
Theorem 1.1 gives a criterion to determine the following $({C}^{\alpha},{C}^{\alpha^{\prime}})$-regularity given by T. C. Dinh, X. N. Ma, and V. A. Nguyen in [DMN17].
Definition 1.5.
For $0<\alpha,\alpha^{\prime}\leq 1$, a compact subset $K$ in a compact Hermitian manifold $X$ is called $({C}^{\alpha},{C}^{\alpha^{\prime}})$-regular if the weighted extremal function $V_{K,\phi}$ of $(K,\phi)$ is $\alpha^{\prime}$-Hölder continuous for any $\alpha$-Hölder continuous weight $\phi$ and each bounded subset in the Hölder space ${C}^{\alpha}(K)$ is sent by the map $\phi\mapsto V_{K,\phi}$ to a bounded subset in ${C}^{\alpha^{\prime}}(X)$.
Theorem 1.6.
Let $K$ be a compact subset of a compact Hermitian manifold $X$. If $K$ has weak local Hölder continuity property, then $K$ is $({C}^{\alpha(K)},{C}^{\alpha^{\prime}(K)})$-regular.
The proof of Theorem 1.6 uses Theorem 1.1 and Theorem 1.2. It also shows that, if $K$ has weak $\mu$-Hölder continuity of order $q$ for constants $\mu\in(0,1]$ and $q>0$, then
$\alpha(K):=\frac{2\mu}{q+2}\quad\text{and}\quad\alpha^{\prime}(K):=\frac{4\mu^{2}}{(q+2)(2\mu+(q+2)^{2})}$
are valid.
Let us go to the research of the Hölder continuity of the extremal functions on $\mathbb{C}^{n}$. The HCP of a general compact subset in $\mathbb{C}^{n}$ (introduced in Section 3) does not imply the local HCP (introduced in Section 3) of the subset . However, if such an compact subset is convex, then the author proves that the $\mu$-HCP of the subset implies the local $\mu$-HCP of it with the order $1$ for each $\mu\in(0,1]$.
Theorem 1.7.
Let $F$ be a compact convex subset of $\mathbb{C}^{n}$ and $\mu\in(0,1]$. If $F$ has $\mu$-HCP, then $F$ has local $\mu$-HCP of order $1$.
Nguyen [Ng24, Corollary 4.6] proved that a compact subset $F$ in $\mathbb{K}^{n}$, where $\mathbb{K}$ is either $\mathbb{R}$ or $\mathbb{C}$, with the geometrical condition in $\mathbb{K}^{n}$ has local $1/2$-HCP of order $n$.
Definition 1.8.
Let $F$ be a compact subset of $\mathbb{K}^{n}$, where $\mathbb{K}$ is either $\mathbb{R}$ or $\mathbb{C}$. $F$ is said to satisfy the geometrical condition if there exists a constant $r_{F}>0$ such that, for each $a\in F$, there exists $a^{\prime}\in F$ letting the convex hull of $\{a\}\cup\bar{\mathbb{B}}_{\mathbb{K}^{n}}(a^{\prime},r_{F})$ be contained in $F$.
Such an $F$ in Definition 1.8 automatically has $1/2$-HCP. Then by Theorem 1.7, we get the following corollary.
Corollary 1.9.
Let $\mathbb{K}$ be either $\mathbb{R}$ or $\mathbb{C}$. Then a convex compact subset of $\mathbb{K}^{n}$ with the nonempty interior in $\mathbb{K}^{n}$ has local $1/2$-HCP of of order $1$.
Organization. In Section 2, we give several basic properties of extremal functions on a compact Hermitian manifold $X$. In Section 3, we introduce the local Hölder continuity property, the weak local Hölder continuity property and the uniform density in capacity of a compact subset in $X$. After the introduction of various Hölder continuity properties, we prove Theorem 1.7. We also study the relation of extremal functions of compact subsets in $\mathbb{C}^{n}$ with uniform densities in capacity of the sets. In Section 4, with the help of Demailly’s estimates, we verify that the global $\mu$-Hölder continuity of the extremal function of a compact subset in $X$ is implied by the $\mu$-Hölder continuity on the set itself. Finally, in Section 5, we prove Theorem 1.1, Theorem 1.2 and Theorem 1.6. Section 6 is an appendix including the lemma explaining the relation between extremal functions on $X$ and locally defined weighted relative extremal functions in $\mathbb{C}^{n}$, which is used in the proof of Theorem 1.2.
Acknowledgement. I would like to thank my advisor Ngoc Cuong Nguyen with my deep gratitude for helping me write this article. This work was supported by the National Research Foundation of Korea (NRF) funded by the Korean government (MSIT) RS-2026-25470686.
2. Basic properties of extremal functions
Throughout this article, $(X,\omega)$ is a compact Hermitian manifold of complex dimension $n$. We impose no additional assumption on the Hermitian metric $\omega$.
The basic properties introduced in this section are the analogues for the extremal functions of the compact sets in $\mathbb{C}^{n}$. For the details of these basic properties, we refer to [Ahn26a]. Note that for subsets $E$ and $O$ of $X$ (resp. of $\mathbb{C}^{n}$), if $O$ is open in $X$ (resp. in $\mathbb{C}^{n}$), then $\sup_{U}V_{E}=\sup_{U}V_{E}^{*}$ (resp. $\sup_{O}L_{E}=\sup_{O}L_{E}^{*}$).
The following proposition by Vu [Vu19, Lemma 2.2, 2.7] gives $\omega$-psh functions. For an open subset $U$ of $X$, $PSH(U,\omega)$ is defined by
$PSH(U,\omega):=\{v:U\to\mathbb{R}\cup\{-\infty\}:v\text{ is quasi-psh},\,\omega+dd^{c}v\geq 0\}.$
Proposition 2.1.
(i) If $U_{1},U_{2}\subset X$ are open with $\overline{U_{1}}\subset U_{2}$, $v_{1}\in PSH(U_{1},\omega)$ and $v_{2}\in PSH(U_{2},\omega)$ with $\limsup_{U_{1}\ni y\to x}v_{1}(y)\leq v_{2}(x)$ for each $x\in\partial U_{1}$, then $v:U_{2}\to[-\infty,\infty)$ defined as $\max\{v_{1},v_{2}\}$ on $U_{1}$ and $v_{2}$ on $U_{2}\setminus U_{1}$ is in $PSH(U_{2},\omega)$.
(ii) If $\{v_{\beta}\}_{\beta\in I}\subset PSH(X,\omega)$ are uniformly bounded above, then $(\sup_{\beta\in I}v_{\beta})^{*}\in PSH(X,\omega)$.
Pluripolar sets and $\omega$-pluripolar sets are also defined on $X$.
Definition 2.2.
Let $E$ be a subset of $X$. $E$ is called pluripolar if for each $p\in E$, there exists open $U\ni p$ and $\phi\in PSH(U)$ such that $\phi\not\equiv$-$\infty$ and $\cap U\subset\{\phi=-\infty\}. EE$ is called $\omega$-pluripolar if $E\subset\{v=-\infty\}$ for some $v\in PSH(X,\omega)$.
We have the following equivalence due to Vu in [Vu19, Theorem 1.1].
Proposition 2.3.
$E\subset X$ is pluripolar if and only if it is $\omega$-pluripolar.
Characterization of pluripolar sets using their extremal functions is possible.
Proposition 2.4.
Let $E$ be a subset of $X$.
  • (i)
    E is pluripolar $\Leftrightarrow$ $\sup_{X}V_{E}^{*}=\infty$ $\Leftrightarrow$ $V_{E}^{*}\equiv\infty$.
  • (ii)
    If E is not pluripolar, then $V_{E}^{*}\in PSH(X,\omega)$ and $V_{E}^{*}|_{int(E)}\equiv 0$.
We also have the following properties as in [GZ17, Proposition 9.19].
Proposition 2.5.
Let $E,F,P$ be subsets of $X$.
  1. (i)
    $E\subset F$ implies $V_{F}\leq V_{E}$.
  2. (ii)
    If $E$ is open, $V_{E}=V^{*}_{E}$.
  3. (iii)
    If $P$ is pluripolar, $V_{E\cup P}^{*}=V_{E}^{*}$.
  4. (iv)
    If $E_{j}\subset X$ is an increasing sequence and $E$ is the limit, $\lim_{j\to\infty}V_{E_{j}}^{*}=V_{E}^{*}.$
  5. (v)
    If compact $K_{j}\subset X$ is a decreasing sequence and $K$ is the limit,
    $\lim_{j\to\infty}V_{K_{j}}=V_{K}$ and $\lim_{j\to\infty}V_{K_{j}}^{*}=V_{K}^{*}$ a.e..
The lower semicontinuity of the extremal functions on $X$ for compact sets holds.
Lemma 2.6.
If $K$ is a compact subset of $X$, then $V_{K}$ is lower semicontinuous.
Corollary 2.7.
If $K$ is a compact subset of $X$, then $V_{K}$ is continuous if and only if $V_{K}^{*}=0$ on $K$.
Remark 2.8.
Corollary  2.7 implies that the continuity of $V_{K}$ for compact $K$ does not depend on the Hermitian metric $\omega$ since for another hermitian metric $\omega^{\prime}$, $\omega^{\prime}\leq c\omega$ for some constant $c>0$, $V_{\omega^{\prime};K}^{*}\leq cV_{\omega;K}^{*}=cV_{K}$. Later, Lemma 4.3 shows that the Hölder continuity of $V_{K}$ and the Hölder exponent are also independent of $\omega$.
The (zero-one) relative extremal function for a set $E$ in $X$ is defined by
(2.1) $h_{E}(z):=\sup\left\{v(z):v\in PSH(X,\omega),\;v\leq 1,\;v|_{E}\leq 0\right\}.$
By Proposition 2.1, $h_{E}^{*}\in PSH(X,\omega)$. Lemma 2.9 proves that $E$ is pluripolar if and only if $h_{E}^{*}\equiv 1$ on the compact Hermitian manifold $X$. For a non-pluripolar compact set $K\subset X$, denote $M_{K}:=\sup_{X}V_{K}^{*}=\sup_{X}V_{K}<\infty$. Then we have
(2.2) $V_{K}^{*}\leq M_{K}h_{K}^{*}.$
The global Bedford-Taylor capacity is comparable (bi-Lipschitz equivalent) to the local Bedford-Taylor capacity. The proof is similar to the compact Kähler case as in [Ko05, page 52-53], [GZ17, Proposition 9.8]. This comparability, Proposition 2.3, [GZ17, Corollary 4.36] and [GZ17, Theorem 4.40] together characterize Borel pluripolar sets and pluripolar sets via the global Bedford-Taylor capacity and the outer global Bedford-Taylor capacity respectively. Characterization of a pluripolar set using its relative extremal function is also possible. The outer global Bedford-Taylor capacity of a subset $E$ in $X$ is defined by
$Cap_{\omega}^{*}(E):=\inf\{Cap_{\omega}(O):O\;\text{open, }E\subset O\subset X\}.$
Lemma 2.9.
Let $\rho_{j}$, $1\leq j\leq N<\infty$, be finitely many smooth strictly psh functions on some holomorphic charts of $X$ properly containing the closures of $U_{j}=\{\rho_{j}<0\}\neq\emptyset$ respectively. Assume for some constant $\delta>0$, $U_{j,\delta}=\{\rho_{j}<-\delta\}$ cover $X$. Then there exist some constants $C_{j}=C_{j}(\rho_{j},X,\omega)>1$, $C_{j}^{\prime}=C_{j}^{\prime}(\rho_{j},X,\omega,\delta)>1$ such that for each $1\leq j_{0}\leq N$ and Borel subset $E\subset X$,
(2.3) $Cap_{\omega}(E\cap U_{j_{0},\delta})$ $\leq C_{j_{0}}Cap(E\cap U_{j_{0},\delta},U_{j_{0}}),$
$Cap_{\omega}(E)\leq\sum_{j=1}^{N}Cap_{\omega}({E\cap U_{j,\delta}})$ $\leq(\max_{1\leq j\leq N}C_{j})\sum_{j=1}^{N}Cap({E\cap U_{j,\delta}},U_{j}),$
(2.4) $Cap(E\cap U_{j_{0},\delta},U_{j_{0}})$ $\leq C_{j_{0}}^{\prime}Cap_{\omega}(E\cap U_{j_{0},\delta})\leq C_{j_{0}}^{\prime}Cap_{\omega}(E),$
$\sum_{j=1}^{N}Cap(E\cap U_{j,\delta},U_{j})$ $\leq(\sum_{j=1}^{N}C_{j}^{\prime})Cap_{\omega}(E).$
Consequently, a Borel subset $P_{1}\subset X$ is pluripolar if and only if $Cap_{\omega}(P_{1})=0$, and a subset $P_{2}\subset X$ is pluripolar if and only if $Cap_{\omega}^{*}(P_{2})=0$ if and only if $h_{P_{2}}^{*}\equiv 1$.
For a non-pluripolar compact subset $K\subset X$, Vu [Vu19, Proposition 2.9] obtained the following estimate of $M_{K}$ in (2.2) using the global Bedford-Taylor capacity.
Lemma 2.10.
There exists a uniform positive constant $A=A(X,\omega)$ depending only on $(X,\omega)$ such that, for every non-pluripolar compact subset $K$ in $X$,
(2.5) $0<\frac{(\int_{X}(dd^{c}V_{K}^{*}+\omega)^{n})^{\frac{1}{n}}}{[Cap_{\omega}(K)]^{\frac{1}{n}}}\leq\max\{1,\sup_{X}V_{K}\},\quad\sup_{X}V_{K}\leq\frac{A}{Cap_{\omega}(K)}.$
For subsets $E_{1}\subset E_{2}$ in $X$ and real-valued functions $\phi_{1}\leq\phi_{2}$ on $X$, there is also monotonicity for weighted extremal functions as
(2.6) $V_{E_{2},\phi_{2}}\leq V_{E_{1},\phi_{1}}\leq V_{E_{1},\phi_{2}}.$
The following lemma tells other basic properties of weighted extremal functions.
Lemma 2.11.
For a compact subset $K$ in $X$ and a bounded function $\phi:X\to\mathbb{R}$,
  • (i)
    $V_{K}^{*}+\inf_{K}\phi\leq V_{K,\phi}^{*}\leq V_{K}^{*}+\sup_{K}\phi.$
  • (ii)
    for the current $\theta:=\omega+dd^{c}\phi$ of bidegree $(1,1)$ and the function $V_{\theta;K}:=\sup\{v\in PSH(X,\theta):v|_{K}\leq 0\}$ where $PSH(X,\theta):=\{v:X\to\mathbb{R}\cup\{-\infty\}:v\text{ is quasi-psh},\,\theta+dd^{c}v\geq 0\}$, $V_{K,\phi}=V_{\omega;K,\phi}=V_{\theta;K}+\phi.$
  • (iii)
    Assume further that $\phi$ is lower semicontinuous. Then $V_{K,\phi}$ is lower semicontinuous. Consequently, for such a function $\phi$, $V_{K,\phi}$ is continuous if and only if $V_{K,\phi}^{*}\leq\phi$ on $K$.
3. Local Hölder continuity property and uniform density in capacity
3.1. Moduli of continuity of extremal functions on complex coordinate spaces
When $E\subset\mathbb{C}^{n}$ is non-pluripolar, the modulus of continuity of $L_{E}$ at $a\in E$ is defined by
$\varpi_{E}(a,\delta):=\sup_{\|z-a\|\leq\delta}L_{E}(z),\quad\delta\in[0,\infty).$
The modulus of continuity of $L_{E}$ on $E$ is defined by
$\varpi_{E}(\delta):=\sup\{\varpi_{E}(a,\delta):a\in E\},\quad\delta\in[0,\infty).$
$L_{E}$ is continuous at $a$ if and only if $\lim_{\delta\to 0^{+}}\varpi_{E}(a,\delta)=0$. Moreover,
(3.1) $|L_{E}(z)-L_{E}(w)|\leq\varpi_{E}(\|z-w\|),\quad z,w\in\mathbb{C}^{n}.$
This can be proved as follows. For each $z^{\prime}\in\mathbb{C}^{n}$ and competitor $u$ for $L_{E}$, $u(\cdot+z^{\prime})-\varpi_{E}(\|z^{\prime}\|)\leq L_{E}(\cdot)$ by definition of $L_{E}$, thus $L_{E}(z+z^{\prime})-\varpi_{E}(\|z^{\prime}\|)\leq L_{E}(z)$.
3.2. Local Hölder continuity property
We introduce the local Hölder continuity property and the weak local Hölder continuity property, which are related to the Hölder continuity of extremal functions on $X$. The term Hölder continuity property is abbreviated as HCP.
Definition 3.1.
Let $q>0,\mu\in(0,1]$ be constants. Let $F$ be a compact subset of $\mathbb{C}^{n}$. $F$ is said to have
  • (i)
    $\mu$-HCP at $a\in F$ if there exists a constant $C>0$ satisfying
    $L_{F}(z)\leq C\delta^{\mu},\quad 0<\delta\leq 1,\,\|z-a\|\leq\delta.$
  • (ii)
    $\mu$-HCP if there exists a constant $C>0$ satisfying
    $L_{F}(z)\leq C\delta^{\mu},\quad 0<\delta\leq 1,\,z\in\bar{\mathbb{B}}(F,\delta).$
  • (iii)
    local $\mu$-HCP of order $q$ at $a\in F$ if there exists a constant $C>0$ satisfying
    $\varpi_{F\cap\bar{\mathbb{B}}(a,r)}(a,\delta)\leq\frac{C\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 1,\,0<r\leq 1.$
  • (iv)
    local $\mu$-HCP of order $q$ if there exists a constant $C>0$ satisfying
    $\varpi_{F\cap\bar{\mathbb{B}}(a,r)}(a,\delta)\leq\frac{C\delta^{\mu}}{r^{q}},\quad a\in F,\,0<\delta\leq 1,\,0<r\leq 1.$
Let $K$ be a compact subset of $X$. $K$ is said to have
  • (v)
    local $\mu$-HCP of order $q$ at $b\in K$ if for each holomorphic coordinate ball $(\Omega,\tau)$ centered at $b$ of radius in $(0,\infty]$, there exists a constant $C>0$ satisfying
    $\varpi_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r)}(\mathbf{0},\delta)\leq\frac{C\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 1,\,0<r\leq 1.$
  • (vi)
    local $\mu$-HCP of order $q$ if for each finite open cover $\{(O_{s},g_{s})\}_{s\in J}$ of $X$ of holomorphic charts and each regular refined open cover $\{(\hat{O}_{s},g_{s})\}$ of $X$ ($\hat{O}_{s}$ is relatively compact in $O_{s}$ for each $s\in J$), there exists a constant $C>0$ such that, for each $(b,s)\in K\times J$ of $b\in K\cap\hat{O}_{s}$,
    $\varpi_{g_{s}(K\cap O_{s})\cap\bar{\mathbb{B}}(g_{s}(b),r)}(g_{s}(b),\delta)\leq\frac{C\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 1,\,0<r\leq 1.$
    Equivalently, $K$ has local $\mu$-HCP of order $q$ if for each finite open cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of $X$ of holomorphic coordinate balls $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)\subset\mathbb{C}^{n}$ with the refined open cover $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$, there exists a constant $C^{\prime}>0$ such that, for each $(b,j)\in K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
    $\varpi_{\tau_{j}(K\cap\Omega_{j})\cap\bar{\mathbb{B}}(\tau_{j}(b),r)}(\tau_{j}(b),\delta)\leq\frac{C^{\prime}\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 1,\,0<r\leq 1.$
  • (vii)
    weak local $\mu$-HCP of order $q$ at $b\in K$ if for some holomorphic coordinate ball $(\Omega,\tau)$ centered at $b$ of radius in $(0,\infty]$, there exists a constant $C>0$ satisfying
    $\varpi_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r)}(\mathbf{0},\delta)\leq\frac{C\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 1,\,0<r\leq 1.$
  • (viii)
    weak local $\mu$-HCP of order $q$ if for some finite open cover $\{(O_{s},g_{s})\}_{s\in J}$ of $X$ of holomorphic charts with some regular refined open cover $\{(\hat{O}_{s},g_{s})\}$ of $X$ ($\hat{O}_{s}$ is relatively compact in $O_{s}$ for each $s\in J$), there exists a constant $C>0$ such that, for each $(b,s)\in K\times J$ of $b\in K\cap\hat{O}_{s}$,
    $\varpi_{g_{s}(K\cap O_{s})\cap\bar{\mathbb{B}}(g_{s}(b),r)}(g_{s}(b),\delta)\leq\frac{C\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 1,\,0<r\leq 1.$
    Equivalently, $K$ has weakly local $\mu$-HCP of order $q$ if for some finite open cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of $X$ of holomorphic coordinate balls of $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)$ with the refined covering $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$, there exists a constant $C^{\prime}>0$ such that, for each $(b,j)\in K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
    $\varpi_{\tau_{j}(K\cap\Omega_{j})\cap\bar{\mathbb{B}}(\tau_{j}(b),r)}(\tau_{j}(b),\delta)\leq\frac{C^{\prime}\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 1,\,0<r\leq 1.$
Local HCP implies HCP for a compact subset in $\mathbb{C}^{n}$ and local HCP implies weak local HCP for a compact subset in $X$. For all the items, checking the inequality for $0<r\leq r_{0}$ for some $r_{0}\in(0,1]$ is enough due to the monotonicity of extremal functions with respect to the set inclusion. For (i), (ii), (iii), (v) and (vii), checking the inequality for $0<\delta\leq\delta_{0}$ for some $\delta_{0}\in(0,1]$ is enough by local boundedness of extremal functions of non-pluripolar sets. For (iv), (vi) and (viii), checking the inequality for $0<\delta\leq\delta_{0}$ for some $\delta_{0}\in(0,1]$ is enough due to inequality (3.1). Also by (3.1), $L_{F}$ is locally $\mu$-Hölder continuous on $\mathbb{C}^{n}$ if and only if $F$ has $\mu$-HCP.
The equivalence in (vi) of this definition is explained as follows. For each $s\in J$, since $\overline{\hat{O}_{s}}$ is compact and is in $O_{s}$, there exists positive integer $N_{s}$ and $N_{s}$ pairs $(b_{s,t},r_{s,t})$ in $X\times(0,\infty)$ for $1\leq t\leq N_{s}$ such that
$\overline{\hat{O}_{s}}\subset\cup_{1\leq t\leq N_{s}}\mathbb{B}(b_{s,t},r_{s,t}),\quad\cup_{1\leq t\leq N_{s}}\mathbb{B}(b_{s,t},5r_{s,t})\subset O_{s}.$
$(\Omega_{s,t},\tau_{s,t}):=(g_{s}^{-1}(\mathbb{B}(b_{s,t},5r_{s,t})),r_{s,t}^{-1}(g_{s}-g_{s}(b_{s,t})),\quad s\in J,\,1\leq t\leq N_{s},$
are holomorphic coordinate balls $\tau_{s,t}(\Omega_{s,t})=\mathbb{B}(\mathbf{0},5)$ with the refined open cover $\{(\tau_{s,t}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{s,t})\}_{s\in J,1\leq t\leq N_{s}}$ of $X$. Let $(b,s)\in K\times J$ be a pair of $b\in K\cap\hat{O}_{s}$. Then there exists $1\leq t\leq N_{s}$ such that $b\in\tau_{s,t}^{-1}(\mathbb{B}(\mathbf{0},1))$. Equivalence follows from
$\varpi_{g_{s}(K\cap O_{s})\cap\bar{\mathbb{B}}(g_{s}(b),r_{s,t}r)}(g_{s}(b),r_{s,t}\delta)=\varpi_{\tau_{s,t}(K\cap\Omega_{s,t})\cap\bar{\mathbb{B}}(\tau_{s,t}(b),r)}(\tau_{s,t}(b),\delta)$
for $0<\delta<\infty$, $0<r<5$. Equivalence in (viii) is also explained by this inequality.
Remark 3.2.
In Definition 3.1, if $K$ has local $\mu$-HCP of order $q$ at $b\in K$, then letting $\delta\to 0^{+}$ in the inequality (v) implies that $K$ is locally $L$-regular at $b$ in the sense of [Ahn26a]. We in [Ahn26a] showed that local $L$-regularity of $K$ at $b$ gives the continuity of $V_{K}$ at $b$. Thus in this case, we know that $V_{K}$ is continuous at $b$.
Moreover, in Definition 3.1, if $K$ has weak local $\mu$-HCP of order $q$ at $b\in K$, then letting $\delta\to 0^{+}$ in the inequality (v) implies that $K$ is weakly locally $L$-regular at $b$ in the sense of [Ahn26a]. We in [Ahn26a] showed that weak local $L$-regularity of $K$ at $b$ gives the continuity of $V_{K}$ at $b$. Accordingly, even in this weaker condition, we also know the continuity of $V_{K}$ at $b$.
However, to guarantee the Hölder continuity of extremal functions on $X$, weak local $L$-regularity and local $L$-regularity are not sufficient, and we need these stronger properties.
Proof of Theorem 1.7.
. By the HCP condition, we have
$L_{F}(z)\leq C\delta^{\mu},\quad 0<\delta\leq 1,\;\|z-a\|\leq\delta,\;a\in F.$
Let $z^{\prime}\in\mathbb{C}^{n}$ be a point in $\mathbb{C}^{n}$ such that $d_{\mathbb{C}^{n}}(z^{\prime},F)>1$. Then there exists $z^{\prime\prime}\in F$ such that $d_{\mathbb{C}^{n}}(z^{\prime},F)=\|z^{\prime}-z^{\prime\prime}\|$. Let $N(z^{\prime}):=\lceil d_{\mathbb{C}^{n}}(z^{\prime},F)\rceil>1$ be the smallest positive integer that is larger than or equal to $d_{\mathbb{C}^{n}}(z^{\prime},F)>1$. Then there exist $N(z^{\prime})+1$ points $\xi_{0}:=z^{\prime}$, $\xi_{1}$, …, $\xi_{N(z^{\prime})}:=z^{\prime\prime}$ in $\mathbb{C}^{n}$ such that
$\|\xi_{j}-\xi_{j-1}\|=\frac{d_{\mathbb{C}^{n}}(z^{\prime},F)}{N(z^{\prime})}\leq 1,\quad 1\leq j\leq N(z^{\prime}).$
Then by the telescoping with the help of (3.1), we have
$L_{F}(z^{\prime})$ $=L_{F}(z^{\prime})-L_{F}(z^{\prime\prime})$
$\leq\sum_{j=1}^{N(z^{\prime})}|L_{F}(\xi_{j})-L_{F}(\xi_{j-1})|$
$\leq\sum_{j=1}^{N(z^{\prime})}\varpi_{F}(1)$
$\leq N(z^{\prime})C\leq 2Cd_{\mathbb{C}^{n}}(z^{\prime},F),\quad d_{\mathbb{C}^{n}}(z^{\prime},F)>1.$
Therefore, we get
$L_{F}(z)\leq 2C\max\{d_{\mathbb{C}^{n}}(z,F),(d_{\mathbb{C}^{n}}(z,F))^{\mu}\},\quad z\in\mathbb{C}^{n}.$
Since $F$ is convex, for each $a\in F$ and $0<r\leq 1$, the affinely transformed set a+(Fa)/ra+(F-a)/r contains F. Since $F$ is compact, there exists a small constant $1>\epsilon>0$ such that $\bar{\mathbb{B}}(a,1/\epsilon)$ contains $F$ for each $a\in F$.
For each $a\in F$ and $0<r\leq 1$, let $T_{a,r\epsilon}:\mathbb{C}^{n}\to\mathbb{C}^{n}$ is the invertible affine transformation given by $T_{a,r\epsilon}(z):=(z-a)/(r\epsilon)+a$. Then we have
$L_{F\cap\bar{\mathbb{B}}(a,r)}(z)=L_{T_{a,r\epsilon}(F)\cap\bar{\mathbb{B}}(a,\frac{1}{\epsilon})}(T_{a,r\epsilon}(z))\leq L_{F}(T_{a,r\epsilon}(z)),\quad z\in\mathbb{C}^{n}.$
Now, we can show the local $\mu$-HCP of order $1$ of $F$. For each $z\in\mathbb{C}^{n}$, there exists $a_{z}\in F$ such that $d_{\mathbb{C}^{n}}(z,F)=\|z-a_{z}\|$. Then we get
$L_{F\cap\bar{\mathbb{B}}(a,r)}(z)$ $\leq L_{F}(T_{a_{z},r\epsilon}(z))$
$\leq 2C\max\{d_{\mathbb{C}^{n}}(T_{a_{z},r\epsilon}(z),F),(d_{\mathbb{C}^{n}}(T_{a_{z},r\epsilon}(z),F))^{\mu}\}$
$\leq 2C\max\{\frac{\|z-a_{z}\|}{r\epsilon},(\frac{\|z-a_{z}\|}{r\epsilon})^{\mu}\}$
$\leq\frac{2C}{\epsilon}\frac{\|z-a_{z}\|^{\mu}}{r},\quad d_{\mathbb{C}^{n}}(z,F)=\|z-a_{z}\|\leq 1.$
Therefore, $F$ has local $\mu$-HCP of order $1$. ∎
3.3. Uniform density in capacity
We define a uniform density in capacity of a compact subset of $X$.
Definition 3.3.
Take a finite open cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of $X$ of holomorphic coordinate balls of $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)\subset\mathbb{C}^{n}$ with the refined cover $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$. We say that a compact subset $K$ of $X$ has a uniform density in capacity of order $q$, if there exist constants $q>0$ and $\varkappa>0$ such that, for each $(a,j)\in K\times\Lambda$ with $a\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$, with the notation $\bar{B}_{j}(a,r):=\tau_{j}^{-1}(\bar{\mathbb{B}}(\tau_{j}(a),r))$, the following inequality holds:
(3.2) $\inf_{0<r\leq 1}\frac{Cap_{\omega}(K\cap\bar{B}_{j}(a,r))}{r^{q}}\geq\varkappa.$
We say that $K$ has a uniform density in capacity of order $q$ at $a\in K$ if (3.2) holds for each $j\in\Lambda$ with $a\in\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$.
Remark 3.4.
Let us make the same settings as Definition 3.3. By the monotonicity of the capacity along set inclusion, checking the inequality (3.2) for $0<r\leq r_{0}$ for some $r_{0}\in(0,1]$ is enough for the uniform density in capacity.
Since the Euclidean distance induced from any coordinate chart and the Riemannian distance $d_{\omega}$ for $(X,\omega)$ are comparable (bi-Lipschitz equivalent) on each compact subset contained in the domain of the coordinate chart, $K$ has a uniform density in capacity of order $q$ if and only if there exist constants $R_{0}>0$, $q^{\prime\prime}=q>0$ and $\varkappa^{\prime\prime}>0$ such that for each $a\in K$,
(3.3) $\inf_{0<R\leq R_{0}}\frac{Cap_{\omega}(K\cap\{p\in X:d_{\omega}(a,p)\leq R\})}{R^{q^{\prime\prime}}}\geq\varkappa^{\prime\prime}.$
Also, $K$ has a uniform density in capacity of order $q$ at $a\in K$ if and only if (3.3) holds. Therefore, the property of a uniform density in capacity and the order $q$ are independent of the choice of an open cover, while $\chi$ depends on the cover.
Let $\tilde{\omega}>0$ be another Hermitian metric on $X$. Since $X$ is compact, $\omega$ and $\tilde{\omega}$ are comparable. In other words, for some constant $c^{\prime}\geq 1$, $(c^{\prime})^{-1}\omega\leq\tilde{\omega}\leq c^{\prime}\omega$. Then
$(c^{\prime})^{-n}Cap_{\omega}(\cdot)\leq Cap_{\omega/c^{\prime}}(\cdot)\leq Cap_{\tilde{\omega}}(\cdot)\leq Cap_{c^{\prime}\omega}(\cdot)\leq(c^{\prime})^{n}Cap_{\omega}(\cdot).$
Plus, $d_{\omega}$ and $d_{\tilde{\omega}}$ are bi-Lipschitz equivalent. Therefore, the property of a uniform density in capacity and the order $q$ are independent of the choice of an Hermitian metric, while $\chi$ depends on the Hermitian metric.
3.4. Relation of extremal functions and uniform density in capacity in the complex coordinate space
Suppose $K\subset X$ is a compact subset contained in a single holomorphic coordinate unit ball in the common-center coordinate $5$-radius ball $(\Omega,\tau)$ with $\tau(\Omega)=\mathbb{B}(\mathbf{0},5)$. Then $\bar{\mathbb{B}}(\tau(K),1)\subset\mathbb{B}(\mathbf{0},2)\Subset\mathbb{B}(\mathbf{0},3)\Subset\mathbb{B}(\mathbf{0},5)$. By inequalities (2.3) and (2.4), $K$ has a uniform density in capacity if and only if for some constants $q_{0}>0$ and $\chi_{0}>0$, for each $(a,r)\in K\times(0,1]$,
(3.4) $\frac{Cap(\tau(K)\cap\bar{\mathbb{B}}(\tau(a),r)\cap\mathbb{B}(\mathbf{0},2),\mathbb{B}(\mathbf{0},3))}{r^{q_{0}}}\geq\varkappa_{0}.$
$\mathbb{C}^{n}$-version of (2.5), [AT84, Theorem 2.1] or [Ko05, Theorem 2.7], tells that inequality (3.4) is equivalent to the existence of constants $A_{1}>0$ and $q_{1}>0$ satisfying
(3.5) $\sup_{\bar{\mathbb{B}}(\mathbf{0},3)}L_{\tau(K)\cap\bar{\mathbb{B}}(\tau(a),r)}^{*}\leq\frac{A_{1}}{r^{q_{1}}},\quad(a,r)\in K\times(0,1].$
Specifically, if (3.4) holds, then (3.5) holds with $q_{1}=q_{0}$ and $A_{1}=\varkappa_{0}^{-1}G(2)$ for some function $G(t)\in(0,\infty)$ only depending on $t\in(0,\infty)$. Conversely, if (3.5) holds, then (3.4) holds with $q_{0}=nq_{1}$ and $\varkappa_{0}=(2\pi/A_{1})^{n}$: for each $(a,r)\in K\times(0,1]$,
(3.6) $\frac{Cap(\tau(K)\cap\bar{\mathbb{B}}(\tau(a),r),\mathbb{B}(\mathbf{0},3))}{r^{nq_{1}}}\geq\frac{\int_{\mathbb{C}^{n}}(dd^{c}\frac{1}{2}\log{(\|z\|^{2}+1)})^{n}}{(A_{1})^{n}}=\frac{(2\pi)^{n}}{(A_{1})^{n}}.$
The following lemma (proved by Nguyen [Ng24, Lemma 3.3]) tells that local HCP implies a uniform density in density in $\mathbb{C}^{n}$.
Lemma 3.5.
Let $K$ be a compact subset of $\mathbb{B}(\mathbf{0},1)\subset\mathbb{C}^{n}$. Suppose $K$ has local $\mu$-HCP of order $q$. Then $K$ has a uniform density in capacity of order $nq$.
Proof.
We can require $K\subset\mathbb{B}(\mathbf{0},1)$. There exist uniform constants $C>0$ such that
$\varpi_{K\cap\bar{\mathbb{B}}(a,r)}(a,\delta)\leq\frac{C\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 4,\quad 0<r\leq 1,\quad a\in K.$
$\sup_{\bar{\mathbb{B}}(\mathbf{0},3)}L_{K\cap\bar{\mathbb{B}}(a,r)}(z)\leq 4^{\mu}C/r^{q},\quad 0<r\leq 1,\quad a\in K.$
$\frac{Cap(K\cap\bar{\mathbb{B}}(a,r),\mathbb{B}(\mathbf{0},3))}{r^{nq}}\geq\frac{(2\pi)^{n}}{(4^{\mu}C)^{n}},\quad 0<r\leq 1,\quad a\in K.$
The last inequality is due to (3.6). ∎
4. Extension of Hölder continuity of extremal functions on compact subsets
4.1. Demailly estimates
The following Demailly estimates in [De94] help us to prove that the Hölder continuity of extremal function $V_{K}$ on $K$ for a compact subset $K$ in $X$ implies the Hölder continuity on $X$. Since $X$ is compact, there exists a constant $R_{\omega}>0$ depending only on $(X,\omega)$ such that, for each function $v\in PSH(X,\omega)\cap L^{\infty}(X)$ and real number $0<\delta\leq R_{\omega}$, the regularization
(4.1) $\Psi_{\delta}v(x):=\frac{1}{\delta^{2n}}\int_{\zeta\in T_{x}^{\prime}X}v({\text{exph}}_{x}(\zeta))\chi\Big(\frac{|\zeta|^{2}_{\omega(x)}}{\delta^{2}}\Big)\,dV_{\omega(x)}(\zeta),\quad x\in X,$
is a smooth function on $X$. Here, $\text{exph}_{x}:T_{x}^{\prime}X\to X$, given in [De94], is the formal holomorphic part of the exponential map ${\exp}_{x}:T_{x}^{\prime}X\to X$ with respect to Chern connection of $(X,\omega)$ on the holomorphic tangent space $T_{x}^{\prime}X$ at $x\in X$. $dV_{\omega(x)}(\zeta):=(dd^{c}\|\zeta\|_{\omega(x)}^{2})^{n}/(2^{n}n!)$ is the Lebesgue measure on $(T_{x}^{\prime}X,\omega(x))$ where
$\omega(x)=\sum_{j,k=1}^{n}\frac{\sqrt{-1}h_{j\bar{k}}}{2}dz^{j}\wedge d\bar{z}^{k},\quad\zeta=\sum_{j=1}^{n}\zeta^{j}\frac{\partial}{\partial z^{j}}\in T_{x}^{\prime}X,$
$\|\zeta\|_{\omega(x)}^{2}=\sum_{j,k=1}^{n}\frac{h_{j\bar{k}}}{2}\zeta^{j}\bar{\zeta}^{k},$
so that $\int_{\|\zeta\|_{\omega(x)}\leq 1}dV_{\omega(x)}(\zeta)=\pi^{n}/n!$ (the volume of a unit ball in $\mathbb{C}^{n}$). $\chi:[0,\infty)\rightarrow[0,\infty)$ is a fixed cut-off function supported on $[0,1]$ with $\chi([0,1))\subset(0,\infty)$ and $\int_{\mathbb{C}^{n}}\chi(\|z\|^{2})\,dV(z)=1$. For convention we define $\Psi_{0}v$ as $v$.
We import the following lemma from [KN19, Lemma 4.1], which is a well-modified version of [BD12, Lemma 1.12]. Let $\|\cdot\|_{X}$ be the supremum norm on $X$ and $USC(X)$ be the set of upper semicontinuous functions from $X$ into $\mathbb{R}\cup\{-\infty\}$.
Lemma 4.1.
There exists a constant $c_{0}>0$ depending only on $(X,\omega)$ such that, for a given $v\in PSH(X,\omega)\cap L^{\infty}(X)$, the functions
(4.2) $USC(X)\ni v_{\delta,c}(x):=\inf_{t\in(0,\delta]}(\Psi_{t}v(x)+c_{1}t^{2}+c_{1}t-c\log\frac{t}{\delta}),\quad x\in X$
defined for $(c,\delta)\in(0,\infty)\times(0,R_{\omega}]$ satisfy the following two properties for constants $c_{1}>0$ and $\delta_{1}\in(0,R_{\omega}]$ depending only on $(X,\omega,\|v\|_{X})$:
  • (i)
    $0,\delta_{1}]\ni t\mapsto\Psi_{t}v+c_{1}t^{2}$ is increasing, and the infimum in (4.2) is attained for each $(x,c,\delta)\in X\times(0,\infty)\times(0,\delta_{1}]$.
  • (ii)
    $\omega+dd^{c}v_{\delta,c}\geq-(c_{0}c+2c_{1}\delta)\,\omega$ for each $(c,\delta)\in(0,\infty)\times(0,\delta_{1}]$.
Note that $c_{1}t^{2}+c_{1}t$ in (4.2) was $c_{1}(t^{2}-\delta^{2})+c_{1}(t-\delta)$ in [KN19, Lemma 4.1]. The difference $c_{1}\delta^{2}+c_{1}\delta$ does not depend on $t$, comes out of the infimum and $dd^{c}(c_{1}\delta^{2}+c_{1}\delta)=0$.
[De94, Remark 4.6] proved that $\psi_{\delta}v$ in (4.1) is uniformly (on $X$) $O(\delta^{2})$-close to
(4.3) $v*\chi_{\delta}(p):=\int_{\bar{\mathbb{B}}(\mathbf{0},1)}v\circ z^{-1}(z(p)+\delta s)\chi(\|s\|^{2})dV(s),\quad p\in z^{-1}(\Omega_{\delta}),$
where $z:U\to\Omega\subset{\mathbb{C}}^{n}$ is a holomorphic coordinate map such that $z(x)=\mathbf{0}$ and $(\partial/\partial z^{j})_{j}$ is an orthonormal basis of $T_{x}^{\prime}X$. $\Omega_{\delta}:=\{y\in\Omega:d(y,\partial\Omega)>\delta\}$. For this $z$,
(4.4) $\check{v}_{\delta}(p):=\frac{1}{V_{2n}(\bar{\mathbb{B}}(\mathbf{0},1))\delta^{2n}}\int_{\bar{\mathbb{B}}(\mathbf{0},\delta)}v\circ z^{-1}(z(p)+y)dV_{2n}(y),\quad p\in z^{-1}(\Omega_{\delta})$
can be compared to $v*\chi_{\delta}$ on $\Omega_{\delta}$ by [KN23, (5.11), (5.12)]: there exists a $\delta$-independent constant $C>0$ such that, for each $\delta>0$ with $\Omega_{\delta}\neq\emptyset$,
(4.5) $v*\chi_{\delta}-v\leq C(\check{v}_{\delta}-v),\quad v_{\delta/2}-v\leq C(v*\chi_{\delta}-v).$
(4.3), (4.5) and [GKZ08, Remark 4.2](or [Z21, Theorem 3.2]) together give
Corollary 4.2.
Let $v\in PSH(X,\omega)\cap L^{\infty}(X)$. If constants $C>0$, $\delta_{0}>0$ and $\mu\in(0,1]$ satisfy $\Psi_{\delta}v-v\leq C\delta^{\mu}$ for $\delta\in(0,\delta_{0}]$, then $v$ is $\mu$-Hölder continuous.
4.2. Hölder continuity of extremal functions
We first prove that the Hölder continuity of $V_{K}$ on $K$ implies the Hölder continuity of $V_{K}$ on $X$.
Lemma 4.3.
Let $K$ be a compact subset of $X$ and $\mu\in(0,1)$. Suppose $V_{K}$ is $\mu$-Hölder continuous on $K$, i.e., there exist constants $C>0$ and $\delta_{0}>0$ satisfying $V_{K}(x)\leq C\delta^{\mu}$ for $d_{\omega}(x,K)\leq\delta\leq\delta_{0}$. Then $V_{K}$ is $\mu$-Hölder continuous on $X$.
Proof.
The assumption implies $V_{K}^{*}=0$ on $K$, so $v:=V_{K}$ is continuous by Corollary 2.7. Consider $\Psi_{t}v$ of bounded $\omega$-psh function $v$ as (4.1). By Lemma 4.1, there exist constants $c_{0}>0$, $c_{1}>0$, $\delta_{1}>0$ such that for each $c\in(0,\infty)$,
$v_{\delta,c}(x)=\inf_{t\in(0,\delta]}\left(\Psi_{t}v(x)+c_{1}t^{2}+c_{1}t-c\log\frac{t}{\delta}\right),\quad 0<\delta\leq R_{\omega}$
are upper semicontinuous functions, $0,\delta_{1}]\ni t\mapsto\Psi_{t}v+c_{1}t^{2}$ increasing and
$\omega+dd^{c}v_{\delta,c}\geq-(c_{0}c+2c_{1}\delta)\omega,\quad\delta\in(0,\delta_{1}],\;c\in(0,\infty).$
We can shrink $\delta_{0}\leq 1$ so that, $\delta_{0}\leq\delta_{1}$ and for each $\delta\in(0,\delta_{0}]$,
$f(\delta):=\frac{1}{c_{0}}-2\frac{c_{1}}{c_{0}}\delta^{1-\mu}>0$
makes $c_{\delta}:=f(\delta)\delta^{\mu}$ satisfy $c_{0}c_{\delta}+2c_{1}\delta=\delta^{\mu}$ and gives
(4.6) $v_{\delta}:=\frac{v_{\delta,c_{\delta}}}{1+\delta^{\mu}}\in PSH(X,\omega).$
Since exph is smooth in both variables, $d_{\zeta}(\text{exph}_{x})(\mathbf{0})=\text{Id}_{T_{x}^{\prime}X}$ and $X$ is compact, there exists a Lipschitz constant $D_{\omega}$ depending only on $X,\omega,R_{\omega}$ such that
$d_{\omega}(x,\text{exph}_{x}(\zeta))\leq D_{\omega}\|\zeta\|_{\omega(x)},\quad x\in X,\;\zeta\in T_{x}^{\prime}X,\;\|\zeta\|_{\omega(x)}\leq R_{\omega}.$
We can shrink $\delta_{0}$ once more so that
$v(x)\leq C\delta^{\mu},\quad d_{\omega}(x,K)\leq\delta\leq D_{\omega}\delta_{0}.$
Then for the constant $c_{2}:=C(D_{\omega})^{\mu}+2c_{1}$, for each $(a,\delta)\in K\times(0,\delta_{0}]$,
(4.7) $(1+\delta^{\mu})v_{\delta}(a)$ $\leq\Psi_{\delta}v(a)+c_{1}\delta^{2}+c_{1}\delta$
$\leq\sup_{d_{\omega}(x,a)\leq D_{\omega}\delta}v(x)+c_{1}\delta^{2}+c_{1}\delta$
$\leq c_{2}\delta^{\mu}.$
$v\geq 0$ makes $v_{\delta}\geq 0$, thus
$v_{\delta}(a)\leq c_{2}\delta^{\mu},\quad(a,\delta)\in K\times(0,\delta_{0}].$
This result, the equation (4.6) and the definition of $v$ together give
(4.8) $v_{\delta}\leq v+c_{2}\delta^{\mu}\quad\text{on }X,\quad\delta\in(0,\delta_{0}].$
Using (4.8) and the argument of [DDGHKZ], we can show Hölder continuity of $v$ as follows. Fix $(x,\delta)\in X\times(0,\delta_{0}]$. By Lemma 4.1, the infimum in the definition of $v_{\delta,c_{\delta}}(x)$ is attained at some $t_{0}=t_{0}(x,\delta)\in(0,\delta]$. By (4.6) and (4.8), we get
$\Psi_{t_{0}}v(x)+c_{1}t_{0}^{2}+c_{1}t_{0}-c_{\delta}\log\frac{t_{0}}{\delta}-(1+\delta^{\mu})v(x)\leq c_{2}\delta^{\mu}(1+\delta^{\mu}).$
From this inequality, for the constant $c_{3}=c_{3}(K):=2c_{2}+\sup_{X}v$, we get
$\Psi_{t_{0}}v(x)+c_{1}t_{0}^{2}-c_{\delta}\log\frac{t_{0}}{\delta}-v(x)\leq c_{3}\delta^{\mu}.$
Since $t\mapsto\Psi_{t}v(x)+c_{1}t^{2}$ is increasing on $0,\delta_{1}]$,
$\Psi_{t_{0}}v(x)+c_{1}t_{0}^{2}\geq(\Psi_{t}v(x)+c_{1}t^{2})|_{t=0}=v(x).$
Accordingly, we get $c_{\delta}\log(\delta/t_{0})\leq c_{3}\delta^{\mu}$. Putting $c_{\delta}=f(\delta)\delta^{\mu}$ to this inequality and then relating to the inequality $f(\delta_{0})\leq f(\delta)$, for $\kappa:=\exp(-c_{3}/{f(\delta_{0})})\in(0,1)$, we get $t_{0}\geq\kappa\delta$. Since $\Psi_{\kappa\delta}v(x)+c_{1}(\kappa\delta)^{2}\leq\Psi_{t_{0}}v(x)+c_{1}t_{0}^{2}$, we get
$\Psi_{\kappa\delta}v(x)-v(x)$ $\leq\Psi_{\kappa\delta}v(x)+c_{1}(\kappa\delta)^{2}-v(x)$
$\leq\Psi_{t_{0}}v(x)+c_{1}t_{0}^{2}-v(x)$
$\leq v_{\delta,c_{\delta}}(x)-v(x)$
$=\delta^{\mu}v_{\delta}(x)+(v_{\delta}(x)-v(x)).$
$\|v_{\delta}\|_{L^{\infty}(X)}=\|v_{\delta}\|_{X}$ for $0<\delta\leq\delta_{0}$} are bounded by a constant $c_{4}=c_{4}(\|v\|_{X},1/\kappa)$ as $\delta/t_{0}(x,\delta)\leq 1/\kappa$. Therefore, (4.8) gives $\Psi_{\kappa\delta}v-v\leq(c_{2}+c_{4})\delta^{\mu}$. In other words,
(4.9) $\Psi_{\delta}v-v\leq(c_{2}+c_{4})\kappa^{-\mu}\delta^{\mu},\quad 0<\delta\leq\kappa\delta_{0}.$
Thus we get the $\mu$-Hölder continuity of $v=V_{K}$ by Corollary 4.2. ∎
Weighted extremal functions on $X$ also have the property similar to Lemma 4.3.
Proposition 4.4.
Let $K\subset X$ be a compact subset, $\mu\in(0,1)$ and $\phi$ be a $\mu$-Hölder continuous real-valued function on $X$. Suppose for some constants $C,\delta_{0}>0$, $V_{K,\phi}(x)\leq\phi(x)+C\delta^{\mu}$ for $d_{\omega}(x,K)\leq\delta\leq\delta_{0}$. Then $V_{K,\phi}$ is $\mu$-Hölder continuous.
Proof.
$v:=V_{K,\phi}$ is continuous by Lemma 2.11-(iii) as $v^{*}|_{K}\leq\phi|_{K}$. Let us use the same notations as in Lemma 4.3. Then equation (4.7) becomes
$(1+\delta^{\mu})v_{\delta}(a)\leq\Psi_{\delta}v(a)+c_{1}\delta^{2}+c_{1}\delta\leq\phi(a)+C^{\prime}\delta^{\mu},\quad(a,\delta)\in K\times(0,\delta_{0}],$
where $C^{\prime}$ is the sum of $C(D_{\omega})^{\mu}+2c_{1}$ and the Hölder coefficient of $\phi$. Thus $v_{\delta}(a)\leq\phi(a)+(\|v_{\delta}\|_{X}+C^{\prime})\delta^{\mu}$. This and the definition of $v$ gives
$v_{\delta}\leq v+(c_{4}+C^{\prime})\delta^{\mu},\quad\delta\in(0,\delta_{0}]$
for $\delta$-independent constant $c_{4}$ defined as in the proof of Lemma 4.3. We have got (4.8). The same proof of Lemma 4.3 gives the $\mu$-Hölder continuity of $v=V_{K,\phi}$. ∎
5. Proofs of the main theorems
5.1. Proof of the first theorem
Proof of Theorem 1.1-(i).
The $\mu$-Hölder continuity of $V_{K}$ at $a\in K$ means $V_{K}(x)\leq C_{a}\delta^{\mu}$ for $d_{\omega}(x,a)\leq\delta\leq 1$ for some constant $C_{a}>0$. This implies $V_{K}^{*}(x)\leq C_{a}\delta^{\mu}$ for $d_{\omega}(x,a)\leq\delta<1$. By monotonicity, we can require $r\leq 1$.
Pick $\psi\in C^{\infty}(\mathbb{R},[0,1])$ supported on [1,1][-1,1] of $\psi(0)=1$ and the chart $(U,f)$ of $X$ with $U\supset\bar{B}(a,r)=f^{-1}(\bar{\mathbb{B}}(f(a),r))$. They induce $\chi_{r}\in C^{\infty}(X,[0,1])$ defined as
$\chi_{r}(x):=\begin{cases}1-\psi(\|f(x)-f(a)\|/r),\quad x\in U,\\ 1,\quad x\in X\setminus\bar{B}(a,r).\end{cases}$
$\chi_{r}(a)=0$ and $\chi_{r}\equiv 1$ on $X\setminus\bar{B}(a,r)$. Its $C^{1}$-norm and $C^{2}$-norm are bounded by
$\|\chi_{r}\|_{C^{1}(X)}\leq c_{1}/r,\quad\|\chi_{r}\|_{C^{2}(X)}\leq c_{2}/r^{2}.$
Here, constants $c_{1}>0$ and $c_{2}>0$ are independent of $a$ and $r$, but they depend on the choice of the holomorphic coordinate chart $(U,f)$. Accordingly, since $\omega>0$, $dd^{c}\chi_{r}\geq-(c_{2}^{\prime}/r^{2})\omega$ for some constant $c_{2}^{\prime}\geq 1$ depending only on $(U,f)$. Then for $\varepsilon_{r}:={r^{2}}/{(2c_{2}^{\prime})}$, the function $\varepsilon_{r}\chi_{r}$ is in $PSH(X,\omega/2)$.
Let $v\in PSH(X,\omega)$, $v\leq 1$ and $v\leq 0$ on $K\cap\bar{B}(a,r)$. By the choice of $\chi_{r}$ and $\varepsilon_{r}$, $\varepsilon_{r}(v-\chi_{r})\leq 0$ on $K$ and $\varepsilon_{r}+1/2\leq 1$. These imply $\varepsilon_{r}(v-\chi_{r})\leq V_{K}$. Accordingly,
$\varepsilon_{r}h^{*}_{K\cap\bar{B}(a,r)}-\varepsilon_{r}\chi_{r}\leq V_{K}^{*}.$
This inequality and $V_{K}^{*}(a)=0=\chi_{r}(a)$ give $h_{K\cap\bar{B}(a,r)}^{*}(a)=0$. Then by Lemma 2.9, $K\cap\bar{B}(a,r)$ is not pluripolar and $Cap_{\omega}(K\cap\bar{B}(a,r))>0$. By (2.5),
$\sup_{X}V_{K\cap\bar{B}(a,r)}^{*}\leq A/{Cap_{\omega}(K\cap\bar{B}(a,r))}<\infty.$
For each $(x,\delta)\in X\times(0,\infty)$ of $d_{\omega}(x,a)\leq\delta<1$, since $\chi_{r}(x)\leq c_{1}\delta/r$, we have
$\varepsilon_{r}h_{K\cap\bar{B}(a,r)}^{*}(x)\leq\varepsilon_{r}\chi_{r}(x)+C_{a}\delta^{\mu}\leq(\frac{c_{1}\varepsilon_{r}}{r}+C_{a})\delta^{\mu}.$
By (2.2), $V_{K\cap\bar{B}(a,r)}^{*}(a)\leq\sup_{X}V_{K\cap\bar{B}(a,r)}^{*}h^{*}_{K\cap\bar{B}(a,r)}(a)$. Then $V_{K\cap\bar{B}(a,r)}^{*}(a)=0$. For the constants $C_{a}$, $c_{1}$ and $c_{2}^{\prime}$ which are independent of $r$, we have
(5.1) $V_{K\cap\bar{B}(a,r)}^{*}(x)\leq(\sup_{X}V_{K\cap\bar{B}(a,r)}^{*})(\frac{c_{1}}{r}+\frac{2c_{2}^{\prime}C_{a}}{r^{2}})\delta^{\mu},\quad d_{\omega}(x,a)\leq\delta<1.$
In other words, $V_{K\cap\bar{B}(a,r)}$ is also $\mu$-Hölder continuous at $a$. ∎
This proof of Theorem 1.1-(i) gives the following estimate. Let $\mu\in(0,1]$.
Corollary 5.1.
Suppose $V_{K}$ is continuous for compact $K\subset X$. Let $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ be a finite cover of $X$ by holomorphic coordinate balls of $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)\subset\mathbb{C}^{n}$ and the refined cover $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$. Denote $\bar{B}_{j}(b,r):=\tau_{j}^{-1}(\bar{\mathbb{B}}(\tau_{j}(b),r))$. Let $a\in K$.
  • (i)
    If $V_{K}$ is $\mu$-Hölder continuous (resp. continuous at $a$) for $\mu\in(0,1]$, then there exist constants $\delta_{0},A^{\prime}>0$ and $r_{0}\in(0,1]$ such that, for each (a,j) (resp. for each $j$) of $a\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
    (5.2) $V_{K\cap\bar{B}_{j}(a,r)}(x)\leq\frac{A^{\prime}}{Cap_{\omega}(K\cap\bar{B}_{j}(a,r))}\frac{\delta^{\mu}}{r^{2}},\quad d_{\omega}(x,a)\leq\delta\leq\delta_{0},\,0<r\leq r_{0}.$
  • (ii)
    If there exist constants $\delta_{0},A^{\prime}>0$, $r_{0}\in(0,1]$ and $\mu\in(0,1)$ such that (5.2) holds for each $(a,j)$ (resp. for each j) of $a\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$, then $V_{K}$ is $\mu$-Hölder continuous (resp. continuous at a).
Proof.
(i) follows from (5.1) and (2.5). By the proof of Theorem 1.1-(i), $A^{\prime}$ depends only on the Hölder coefficient of $V_{K}$ and the finite open cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$.
Showing the pointwise Hölder continuity of $V_{K}$ in (ii) is obvious by monotonicity of extremal functions. We show the global Hölder continuity of $V_{K}$ in (ii).
$X$ is a compact metric space with Riemannian distance $d_{\omega}$. Thus by Lebesgue number lemma, some $\eta\in(0,r_{0}]$ can be a Lebesgue number for the refined finite open cover $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$. By Remark 3.4, there exist uniform positive constants $0<l_{1}\leq 1\leq l_{2}$ depending only on $X,\omega,\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ such that, for each $j\in\Lambda$ and $p,p^{\prime}\in\tau^{-1}_{j}(\bar{\mathbb{B}}(\mathbf{0},4))$,
$l_{1}\|\tau_{j}(p)-\tau_{j}(p^{\prime})\|\leq d_{\omega}(p,p^{\prime})\leq l_{2}\|\tau_{j}(p)-\tau_{j}(p^{\prime})\|.$
Obviously, we can shrink $\eta$ to satisfy $\eta\leq l_{1}r_{0}$. Again, since $X$ is compact, there exists a finite open cover $\{D_{\lambda}\}$ of $X$, where each $D_{\lambda}$ has $d_{\omega}$-diameter less than $\eta$.
Let $d_{\omega}(x,K)\leq\delta\leq\delta_{0}$. There exists $a\in K$ of $d_{\omega}(x,a)=d_{\omega}(x,K)$. Then $a\in D_{\lambda_{a}}$ for some $\lambda_{a}$, $D_{\lambda_{a}}\subset\tau_{j_{a}}^{-1}(\mathbb{B}(\mathbf{0},1))$ for some $j_{a}\in\Lambda$. By the bi-Lipschitz equivalence of the distances,
$\bar{B}_{j_{a}}(a,\frac{\eta}{l_{1}})\supset O_{a}:=\{p\in X:d_{\omega}(a,p)\leq\eta\}\supset D_{\lambda_{a}}.$
Whenever $K\cap D_{\lambda}\neq\emptyset$, some $b\in K\cap D_{\lambda}$ and $j_{b}\in\Lambda$ with $b\in\tau_{j_{b}}^{-1}(\mathbb{B}(\mathbf{0},1))$ exist to satisfy (5.2). For some small $r_{b}\in(0,r_{0}]$, $\bar{B}_{j_{b}}(b,r_{b})\subset D_{\lambda}$. By (5.2), $K\cap\bar{B}_{j_{b}}(b,r_{b})$ is non-pluripolar, $Cap_{\omega}(K\cap\bar{B}_{j_{b}}(b,r_{b}))>0$, so $Cap_{\omega}(K\cap D_{\lambda})>0$. Therefore,
$Cap_{\omega}(K\cap\bar{B}_{j_{a}}(a,\frac{\eta}{l_{1}}))$ $\geq Cap_{\omega}(K\cap D_{\lambda_{a}})$
$\geq\min_{\lambda:K\cap D_{\lambda}\neq\emptyset}Cap_{\omega}(K\cap D_{\lambda})=:c$
$>0.$
Then the inclusion $K\supset K\cap\bar{B}_{j_{a}}(a,\eta/l_{1})$ and (5.2) with $\eta/l_{1}\leq r_{0}$ gives
$V_{K}(x)\leq V_{K\cap\bar{B}_{j_{a}}(a,\frac{\eta}{l_{1}})}(x)\leq\frac{\tilde{A}}{c}(\frac{\eta}{l_{1}})^{-2}\delta^{\mu}.$
This inequality and Lemma 4.3 prove the $\mu$-Hölder continuity of $V_{K}$. ∎
With this Corollary, we can prove Theorem 1.1-(ii).
Proof of Theorem 1.1-(ii).
We want to show the Hölder continuity of $V:=V_{K,\phi}$. By taking the minimum of Hölder exponents of $V_{K}$ and $\phi$, we can assume that $V_{K}$ and $\phi$ are Hölder continuous with the same Hölder exponent $0<\mu\leq 1$.
Let $b\in K$ and $\varepsilon^{\prime}>0$. $\phi|_{\bar{B}(b,r)}\leq\phi(b)+\varepsilon^{\prime}$ for small $r>0$. By (2.6),
$V\leq V_{K\cap\bar{B}(b,r),\phi}\leq V_{K\cap\bar{B}(b,r),\phi(b)+\varepsilon^{\prime}}=\phi(b)+\varepsilon^{\prime}+V_{K\cap\bar{B}(b,r)}.$
By Theorem 1.1-(i), $V_{K\cap\bar{B}(b,r)}^{*}(b)=0$. Then $V^{*}(b)\leq\phi(b)+\varepsilon^{\prime}$. $\varepsilon^{\prime}\to 0^{+}$ gives $V^{*}(b)\leq\phi(b)$. Accordingly, $^{*}|_{K}\leq\phi|_{K}, VV$ is continuous by Lemma 2.11. To prove the Hölder continuity of $V$, by Proposition 4.4, it is enough to show that there exist constants $C>0$, $\mu^{\prime}>0$ and $\delta_{0}>0$ satisfying
$V(x)\leq\phi(x)+C\delta^{\mu^{\prime}},\quad d_{\omega}(x,K)\leq\delta\leq\delta_{0}.$
Since $X$ is compact, it has a finite cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of holomorphic balls $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)\subset\mathbb{C}^{n}$ with the refined cover $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$. Denote $\bar{B}_{j}(b,r):=\tau_{j}^{-1}(\bar{\mathbb{B}}(\tau_{j}(b),r))$. By Remark 3.4, there exist positive constants $0<l_{1}\leq 1\leq l_{2}$ such that for each $j\in\Lambda$ and $p,p^{\prime}\in\tau^{-1}_{j}(\bar{\mathbb{B}}(\mathbf{0},4))$,
$l_{1}\|\tau_{j}(p)-\tau_{j}(p^{\prime})\|\leq{\rm dist}(p,p^{\prime})\leq l_{2}\|\tau_{j}(p)-\tau_{j}(p^{\prime})\|.$
Set $\delta_{1}:=(l_{2})^{-1}\leq 1$. Let $x\in X$ with $d_{\omega}(x,K)\leq\delta\leq\delta_{1}$. Pick $b_{x}\in K$ of $d_{\omega}(x,b_{x})=d_{\omega}(x,K)$. Then $b_{x}\in\tau_{j(b_{x})}^{-1}(\mathbb{B}(\mathbf{0},1))$ for some $j(b_{x})\in\Lambda$. Let $r\in(0,\delta_{1}]$ be determined later. By Hölder continuity, $|\phi(w)-\phi(b_{x})|\leq c_{0}(l_{2}r)^{\mu}$ for each $w\in\bar{B}_{j(b_{x})}(b_{x},r)$, where $c_{0}$ is a Hölder coefficient of $\phi$. By the monotonicity (2.6),
(5.3) $V(x)\leq V_{K\cap\bar{B}_{j(b_{x})}(b_{x},r),\phi}(x)$ $\leq V_{K\cap\bar{B}_{j(b_{x})}(b_{x},r),\phi(b_{x})+c_{0}l_{2}^{\mu}r^{\mu}}(x)$
$=\phi(b_{x})+c_{0}l_{2}^{\mu}r^{\mu}+V_{K\cap\bar{B}_{j(b_{x})}(b_{x},r)}(x)$
$\leq\phi(x)+c_{0}\delta^{\mu}+c_{0}l_{2}^{\mu}r^{\mu}+V_{K\cap\bar{B}_{j(b_{x})}(b_{x},r)}(x).$
Corollary 5.1-(i) and a uniform density in capacity of $K$ give constants $0<\delta_{2}\leq 1$, $0<r_{2}<1$, $0<A^{\prime}$, $0<\varkappa$ and $0<q$ which are independent of $b_{x}$ and $j(b_{x})$ such that, for each $(y,\delta,R)\in X\times(0,\infty)^{2}$ of $d_{\omega}(y,b_{x})\leq\delta\leq\delta_{2}$ and $0<R\leq r_{2}$,
(5.4) $V_{K\cap\bar{B}_{j(b_{x})}(b_{x},R)}(y)\leq\frac{A^{\prime}}{Cap_{\omega}(K\cap\bar{B}_{j(b_{x})}(b_{x},R))}\frac{\delta^{\mu}}{R^{2}}\leq\frac{A^{\prime}\,\delta^{\mu}}{\varkappa\,R^{q+2}}.$
$(0,\infty)\ni s\mapsto\min\{s\mu,\mu-s(q+2)\}$ is maximized at $s_{0}:={\mu}/{(\mu+q+2)}\in(0,1)$ with the maximum $\mu^{\prime}:={\mu^{2}}/{(\mu+q+2)}<\mu$. Set $\delta_{0}:=\min\{\delta_{1},\delta_{2},r_{2}\}$. Whenever $d_{\omega}(z,K)\leq\delta\leq\delta_{0}$, there exists $b_{z}\in K$ of $d_{\omega}(z,b_{z})=d_{\omega}(z,K)\leq\delta\leq\min\{\delta_{1},\delta_{2}\}$, and $r=R(\delta):=\delta^{s_{0}}\leq\delta\leq\delta_{0}\leq\min\{\delta_{1},r_{2}\}$. Therefore, by (5.3) and (5.4), for the constant $C:=c_{0}+c_{0}l_{2}^{\mu}+{A^{\prime}}/{\varkappa}>0$, we have
$V(z)\leq\phi(z)+C\delta^{\mu^{\prime}},\quad d_{\omega}(z,K)\leq\delta\leq\delta_{0}.$
Then by Proposition 4.4, $V=V_{K,\phi}$ is Hölder continuous on $X$. ∎
Remark 5.2.
The proof of Theorem 1.1-(ii) demonstrated that for a compact $K$ having a uniform density in capacity with $q$ as in (3.2), $\mu$-Hölder continuities of the extremal function $V_{K}$ of $K$ and $\phi$ implies the $\mu^{\prime}$-Hölder continuity of the weighted extremal function $V_{K,\phi}$, where $\mu^{\prime}=\mu^{2}/(\mu+2+q)<\mu$.
Remark 5.3.
Let $\phi$ be a smooth real-valued function on $X$. Then there exists a constant $c>0$ such that $\tilde{\omega}:=dd^{c}\phi+c\omega>0$ becomes a Hermitian metric. Assume $V_{K,\phi/c}$ is $\nu$-Hölder continuous and $K$ has a uniform density in capacity of order $q$ for some $\nu\in(0,1]$ and $q>0$. By Remark 3.4, $K$ also has a uniform density in capacity of order $q$ with respect to $\tilde{\omega}$. Since $V_{K,\phi/c}=c^{-1}(V_{\tilde{\omega};K}+\phi)$ is $\nu$-Hölder continuous, the function $V_{\tilde{\omega};K}$ is also $\nu$-Hölder continuous. Thus $V_{\tilde{\omega};K}$ is $\nu$-Hölder continuous. For $\nu^{\prime}:=\nu^{2}/(\nu+2+q)$, by Theorem 1.1-(ii), $V_{\tilde{\omega};K,-\phi}$ is $\nu^{\prime}$-Hölder continuous. Therefore, the unweighted extremal function $V_{K}=c^{-1}(V_{\tilde{\omega};K,-\phi}+\phi)$ of $K$ is $\nu^{\prime}$-Hölder continuous.
5.2. Proof of the second theorem
Finally, we prove Theorem 1.2. We first show the equivalence of (i), (ii) and (iii) in Theorem 1.2, where the implication (ii) $\Rightarrow$ (iii) is obvious from the definitions.
Proof of (iii) $\Rightarrow$ (i) in Theorem 1.2.
Assume $K$ has weak local $\mu$-HCP of order $q$ of $K$ at $a\in K$ for constants $0<\mu\leq 1$ and $q>0$. Then there is a holomorphic coordinate ball $(\Omega,\tau)$ centered at $a$ and a constant $C>0$ satisfying
$\varpi_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r)}(\mathbf{0},\delta)\leq\frac{C\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 4,\;0<r\leq 1.$
In this inequality, for $\delta=4$, we have
(5.5) $\sup_{\bar{\mathbb{B}}(\mathbf{0},3)}L^{*}_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r)}\leq\frac{4^{\mu}C}{r^{q}},\quad 0<r\leq 1.$
Here, the range $\delta\in(0,4]$ is available by (3.1). Set $C^{\prime}:=4^{\mu}C$. (3.6) gives
$\frac{Cap(\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r),\mathbb{B}(\mathbf{0},3))}{r^{nq}}\geq\frac{(2\pi)^{n}}{(C^{\prime})^{n}},\quad 0<r\leq 1.$
There exists $R\in(0,\infty]$ such that $\tau(\Omega)=\mathbb{B}(\mathbf{0},R)$). Let us define $R_{1}:=\min\{3,R/2\}$ and $R_{2}:=\min\{2,R_{1}/2\}$. Then we have
$\frac{Cap((\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r))\cap\mathbb{B}(\mathbf{0},R_{2}),\mathbb{B}(\mathbf{0},R_{1}))}{r^{nq}}\geq\frac{(2\pi)^{n}}{(C^{\prime})^{n}},\quad 0<r\leq\frac{R_{2}}{2}.$
Since $0<R_{2}<R_{1}<R$, (2.4) gives some constant $C^{\prime\prime}>0$ satisfying
$\frac{Cap_{\omega}(K\cap\tau^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)))}{r^{nq}}\geq C^{\prime\prime},\quad 0<r\leq\frac{R_{2}}{2}.$
This implies (3.3) by bi-Lipschitz equivalence of the Riemannian distance $d_{\omega}$ and the Euclidean distance on a compact set in a coordinate chart. Therefore, $K$ has a uniform density in capacity of order $nq$ at $a$.
To finish the proof, we want to show the Hölder continuity of $V_{K}$ at $a\in K$. There exists a bounded non-negative smooth strictly psh function of the form $\rho_{2}(\cdot)=D\|\tau(\cdot)\|^{2}$ for some constant $D>0$, on $\tau^{-1}(\mathbb{B}(\mathbf{0},R_{1}))$, satisfying $dd^{c}\rho_{2}\geq\omega$, $\rho_{2}(a)=0$. By Lemma 6.1, for some constant $M\in(0,\infty)$,
$(V_{K\cap\tau^{-1}(\bar{\mathbb{B}}(\mathbf{0},R_{2}))}+\rho_{2})\circ\tau^{-1}\leq Mu_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},R_{2}),\rho_{2}\circ\tau^{-1};\mathbb{B}(\mathbf{0},R_{1})}\quad\text{on }\mathbb{B}(\mathbf{0},R_{1}).$
The inequality (6.3), comparability (6.2) and inequality (5.5) yield that for each $0<r\leq\min\{1,R_{2}\}$ and $\|z\|\leq\delta\leq\min\{1,R_{2}\}$, as $\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r)\subset\bar{\mathbb{B}}(\mathbf{0},R_{2})$,
$u_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r),\rho_{2}\circ\tau^{-1};\mathbb{B}(\mathbf{0},R_{1})}(z)$
$\leq(1+\|\rho_{2}\|_{\tau^{-1}(\mathbb{B}(\mathbf{0},R_{1}))})u_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r);\mathbb{B}(\mathbf{0},R_{1})}(z)+\sup_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r)}\rho_{2}\circ\tau^{-1}$
$\leq\frac{1+\|\rho_{2}\|_{\tau^{-1}(\mathbb{B}(\mathbf{0},R_{1}))}}{\inf_{\partial\mathbb{B}(\mathbf{0},R_{1})}L_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r)}}L_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r)}(z)+\sup_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r)}\rho_{2}\circ\tau^{-1}$
$\leq\frac{1+\|\rho_{2}\|_{\tau^{-1}(\mathbb{B}(\mathbf{0},R_{1}))}}{\inf_{\partial\mathbb{B}(\mathbf{0},R_{1})}L_{\bar{\mathbb{B}}(\mathbf{0},R_{2})}}\frac{C\delta^{\mu}}{r^{q}}+Dr^{2}$
$=\frac{1+DR_{1}^{2}}{\log{R_{1}}-\log{R_{2}}}\frac{C\delta^{\mu}}{r^{q}}+Dr^{2}.$
The following function optimizes the exponent. At $t_{0}:=\mu/{(q+2)}\in(0,1)$,
$(0,\infty)\ni t\mapsto\min\{\mu-qt,2t\}$
attains its maximum $\mu^{\prime}:={2\mu}/{(q+2)}<\mu$. Using this fact, for each $\|z\|\leq\delta\leq(\min\{1,R_{2}\})^{1/t_{0}}$, choosing $r(\delta):=\delta^{t_{0}}\leq\min\{1,R_{2}\}$, we get
$u_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},\delta^{t_{0}})),\rho_{2}\circ\tau^{-1};\mathbb{B}(\mathbf{0},R_{1})}(z)\leq D^{\prime}\delta^{\mu^{\prime}},\quad D^{\prime}=\frac{(1+DR_{1}^{2})C}{\log{R_{1}}-\log{R_{2}}}+D.$
Remember the setting $\rho_{2}\geq 0$. Then for each $\|z\|\leq\delta\leq(\min\{1,R_{2}\})^{1/t_{0}}$,
$V_{K}\circ\tau^{-1}(z)$ $\leq(V_{K\cap\tau^{-1}(\bar{\mathbb{B}}(\mathbf{0},R_{2}))}+\rho_{2})\circ\tau^{-1}(z)$
$\leq Mu_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},R_{2}),\rho_{2}\circ\tau^{-1};\mathbb{B}(\mathbf{0},R_{1})}(z)$
$\leq Mu_{\tau(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},\delta^{t_{0}})),\rho_{2}\circ\tau^{-1};\mathbb{B}(\mathbf{0},R_{1})}(z)\leq MD^{\prime}\delta^{\mu^{\prime}}.$
Therefore, this inequality and equivalence of ($d_{\omega}$, $(\tau\times\tau)^{*}d_{\mathbb{C}^{n}}$) on $\tau^{-1}(\bar{\mathbb{B}}(\mathbf{0},R_{2}))$ tell that $V_{K}$ is $\mu^{\prime}$-Hölder continuous at the point $\tau^{-1}(\mathbf{0})=a\in K$.∎
Proof of (i) $\Rightarrow$ (ii) in Theorem 1.2.
Suppose there exist constants $0<\nu\leq 1$ and $0<q_{0}$ such that $V_{K}$ is $\nu$-Hölder continuous at $a\in K$ and $K$ has a uniform density in capacity of order $q_{0}$ at $a$. We want to the local HCP of $K$ at $a$. Fix a holomorphic coordinate ball $(\Omega_{0},\tau_{0})$ of radius $R_{0}\in(0,\infty]$ centered at $a$. Take $R_{3}:=\min\{1,R_{0}/2\}$. Since $K$ has a uniform density in capacity at $a$ of order $q_{0}$, there exists a constant $\varkappa_{0}>0$ such that
$\frac{Cap_{\omega}(K\cap\tau_{0}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r))}{r^{q_{0}}}\geq\varkappa_{0},\quad 0<r\leq\frac{R_{3}}{2}.$
By (2.3), there exists a bi-Lipschitz constant $C_{0}>0$ such that
$Cap_{\omega}(\cdot\cap\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},\frac{2R_{3}}{3})))\leq C_{0}Cap(\cdot\cap\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},\frac{2R_{3}}{3})),\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},R_{3}))).$
Then for each $0<r\leq R_{3}/2$, since $R_{3}/2<2R_{3}/3$, we have
$Cap_{\omega}(K\cap\tau_{0}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)))\leq C_{0}Cap(K\cap\tau_{0}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)),\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},R_{3}))),$
$\frac{Cap((K\cap\tau_{0}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r))\cap\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},\frac{2R_{3}}{3})),\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},R_{3})))}{r^{q_{0}}}\geq\frac{\varkappa_{0}}{C_{0}},$
$\frac{Cap((\tau_{0}(K\cap\Omega_{0})\cap\bar{\mathbb{B}}(\mathbf{0},r))\cap\mathbb{B}(\mathbf{0},\frac{2R_{3}}{3}),\mathbb{B}(\mathbf{0},R_{3}))}{r^{q_{0}}}\geq\frac{\varkappa_{0}}{C_{0}}.$
Since $Cap(cE,cO)=Cap(E,O)$, for each $0<r\leq R_{3}/2$,
$\frac{Cap(\frac{3}{R_{3}}(\tau_{0}(K\cap\Omega_{0})\cap\bar{\mathbb{B}}(\mathbf{0},r))\cap\mathbb{B}(\mathbf{0},2),\mathbb{B}(\mathbf{0},3))}{r^{q_{0}}}\geq\frac{\varkappa_{0}}{C_{0}}.$
By (3.5), there exists a constant $C_{1}>0$ such that, for each $0<r\leq R_{3}/2$,
(5.6) $\sup_{\bar{\mathbb{B}}(\mathbf{0},R_{3})}L_{(\tau_{0}(K\cap\Omega_{0})\cap\bar{\mathbb{B}}(\mathbf{0},r))}^{*}=\sup_{\bar{\mathbb{B}}(\mathbf{0},3)}L_{\frac{3}{R_{3}}(\tau_{0}(K\cap\Omega_{0})\cap\bar{\mathbb{B}}(\mathbf{0},r))}^{*}\leq\frac{C_{1}}{r^{q_{0}}}.$
Since $V_{K}$ is $\nu$-Hölder continuous at $a$ and $K$ has a uniform density in capacity at $a$ of order $q_{0}$, Corollary 5.1-(i) guarantees a constant $A^{\prime}>0$ satisfying
$V_{K\cap\tau_{0}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r))}(x)\leq\frac{A^{\prime}}{\varkappa_{0}}\frac{\delta^{\nu}}{r^{q_{0}+2}},\quad d_{\omega}(x,a)\leq\delta\leq 1,\;0<r\leq\frac{R_{3}}{2}.$
Since $R_{3}<R_{0}$, there exists a bounded non-negative smooth strictly psh function $\rho_{1}$ on $\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},R_{3}))$ of the form $\rho_{1}(\cdot)=d\|\tau_{0}(\cdot)\|^{2}$ for a constant $d>0$ such that
$\|\rho_{1}\|_{\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},R_{3}))}<\infty,\quad\rho_{1}(a)=0,\quad dd^{c}\rho_{1}\leq\omega,\quad\liminf_{x\to\partial(\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},R_{3})))}\rho_{1}(x)>0.$
There exists a bi-Lipschitz constant $l>1$ such that $d_{\omega}$ is dominated by $l\times((\tau_{0}\times\tau_{0})^{*}d_{\mathbb{C}^{n}})$ on $\tau_{0}^{-1}(\bar{\mathbb{B}}(\mathbf{0},R_{3}))$. The comparability (6.2), inequality (6.3) and Lemma 6.1 in turn give that, for each $\|z\|\leq\delta\leq R_{3}/l$ and $0<r\leq R_{3}/2$, with
$m:=\frac{\liminf_{x\to\partial(\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},R_{3})))}\rho_{1}(x)}{1+{\|\rho_{1}\|}_{\tau_{0}^{-1}(\mathbb{B}(\mathbf{0},R_{3}))}}=\frac{R_{3}^{2}d}{1+R_{3}^{2}d}\quad\text{and}$
$c(r):=\sup_{\mathbb{B}(\mathbf{0},R_{3})}L_{\tau_{0}(K\cap\Omega_{0})\cap\bar{\mathbb{B}}(\mathbf{0},r)}\,$,
we have the following chain of inequalities:
$L_{\tau_{0}(K\cap\Omega)\cap\bar{\mathbb{B}}(\mathbf{0},r))}(z)$ $\leq c(r)u_{\tau_{0}(K\cap\Omega_{0})\cap\bar{\mathbb{B}}(\mathbf{0},r);\mathbb{B}(\mathbf{0},R_{3})}(z)$
$\leq c(r)(u_{\tau_{0}(K\cap\Omega_{0})\cap\bar{\mathbb{B}}(\mathbf{0},r),\rho_{1};\mathbb{B}(\mathbf{0},R_{3})}(z)-\inf_{\mathbb{B}(\mathbf{0},R_{3})}\rho_{1})$
$=c(r)u_{\tau_{0}(K\cap\Omega_{0})\cap\bar{\mathbb{B}}(\mathbf{0},r),\rho_{1};\mathbb{B}(\mathbf{0},R_{3})}(z)$
$\leq\frac{c(r)}{m}(V_{K\cap\tau_{0}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r))}(\tau_{0}^{-1}(z))+\rho_{1}(\tau_{0}^{-1}(z)))$
$\leq\frac{c(r)}{m}(\frac{A^{\prime}(l\delta)^{\nu}}{\varkappa_{0}r^{q_{0}+2}}+d\delta^{2})$
$\leq\frac{1+R_{3}^{2}d}{R_{3}^{2}d}C_{1}(\frac{A^{\prime}}{\varkappa_{0}}\frac{l^{\nu}}{r^{2q_{0}+2}}+\frac{d}{r^{q_{0}}})\delta^{\nu}$
$\leq\frac{1+R_{3}^{2}d}{R_{3}^{2}d}C_{1}(\frac{A^{\prime}}{\varkappa_{0}}l^{\nu}+d)\frac{\delta^{\nu}}{r^{2q_{0}+2}}.$
In this chain of inequalities, (5.6) is utilized to dominate $c(r)$. This shows that $K$ has local $\nu$-Hölder continuity property at $a\in K$ of order $2q_{0}+2$. ∎
We have shown the equivalence (i), (ii) and (iii) in Theorem 1.2. The proof of the equivalence (iv), (v) and (vi) in Theorem 1.2 uses the proof of the equivalence (i), (ii) and (iii), with additional uniform controls for global properties. Again, the implication (v) $\Rightarrow$ (vi) in this theorem is obvious from the definitions.
Proof of (vi) $\Rightarrow$ (iv) in Theorem 1.2.
Assume that there exist constants $0<\mu\leq 1$ and $q>0$ such that $K$ has weak local $\mu$-HCP of order $q$ with respect to some finite cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of $X$ of holomorphic coordinate balls $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)$ with its refined cover $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$. We want to show a uniform density in capacity of $K$ and the Hölder continuity of $V_{K}$. Write $\tau_{j}^{-1}(\bar{\mathbb{B}}(\tau_{j}(b),r))=:\bar{B}_{j}(b,r)$ and $\tau_{j}^{-1}(\mathbb{B}(\tau_{j}(b),r))=:B_{j}(b,r)$. For each $j\in\Lambda$, let $b_{j}\in X$ be the point of $\tau_{j}(b_{j})=\mathbf{0}$. There exists a constant $C>0$ such that, for each $(b,j)\in K\times\Lambda$ of $b\in K\cap B_{j}(b_{j},1)=K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
$\varpi_{\tau_{j}(K\cap\bar{B}_{j}(b,r))}(\tau_{j}(b),\delta)\leq\frac{C\delta^{\mu}}{r^{q}},\quad 0<\delta\leq 5,\,0<r\leq 1.$
Here, $\delta\in(0,5]$ instead of $\delta\in(0,1]$ is possible due to (3.1). By Remark 3.4, there exist bi-Lipschitz constants $0<l_{1}\leq 1\leq l_{2}$ such that for each $j\in\Lambda$ and $p,p^{\prime}\in\tau_{j}^{-1}(\bar{\mathbb{B}}(\mathbf{0},4))$,
$l_{1}\|\tau_{j}(p)-\tau_{j}(p^{\prime})\|\leq d_{\omega}(p,p^{\prime})\leq l_{2}\|\tau_{j}(p)-\tau_{j}(p^{\prime})\|.$
Also, there exist uniform constants $d,D>0$ such that, for each pair $(p,j)\in X\times\lambda$ with $p\in\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
$\rho_{1,p,j}(\cdot):=d\|\tau_{j}(\cdot)-\tau_{j}(p)\|^{2},\quad\rho_{2,p,j}(\cdot):=D\|\tau_{j}(\cdot)-\tau_{j}(p)\|^{2}$
are non-negative bounded smooth strictly psh functions on $B_{j}(p,3)$ satisfying
$dd^{c}\rho_{1,p,j}\leq\omega\leq dd^{c}\rho_{2,p,j},\quad\rho_{1,p,j}(p)=0=\rho_{2,p,j}(p),$
$\liminf_{x\to\partial B_{j}(p,3)}\rho_{1,p,j}(x)>0.$
The weak local HCP tells, for each $(b,j)\in\ K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
(5.7) $L_{\tau_{j}(K\cap\bar{B}_{j}(b,r))}(z)\leq\frac{C\delta^{\mu}}{r^{q}},\quad\|z-\tau_{j}(b)\|\leq\delta\leq 5,0<r\leq 1.$
Then for $\delta=5$ and $C^{\prime}:=5^{\mu}C$, for each $(b,j)\in\ K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
$\sup_{\bar{\mathbb{B}}(\mathbf{0},3)}L_{\tau_{j}(K\cap\bar{B}_{j}(b,r))}^{*}\leq\frac{C^{\prime}}{r^{q}},\quad 0<r\leq 1.$
By the inequality (3.6), for each $(b,j)\in\ K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
$\frac{Cap(\tau_{j}(K\cap\bar{B}_{j}(b,r))\cap\mathbb{B}(\mathbf{0},2),\mathbb{B}(\mathbf{0},3))}{r^{nq}}\geq\frac{(2\pi)^{n}}{(C^{\prime})^{n}},\quad 0<r\leq 1,$
$\frac{Cap((K\cap\bar{B}_{j}(b,r))\cap B_{j}(b_{j},2),B_{j}(b_{j},3))}{r^{nq}}\geq\frac{(2\pi)^{n}}{(C^{\prime})^{n}},\quad 0<r\leq 1,$
$\frac{Cap(\{p\in K:d_{\omega}(b,p)\leq l_{2}r\}\cap B_{j}(b_{j},2),B_{j}(b_{j},3))}{r^{nq}}\geq\frac{(2\pi)^{n}}{(C^{\prime})^{n}},\quad 0<r\leq 1,$
as $\{p\in X:d_{\omega}(b,p)\leq l_{2}r\}\supset\bar{B}_{j}(b,r)$. Thus for each $b\in K$,
$\frac{\sum_{i\in\Lambda}Cap(\{p\in K:d_{\omega}(b,p)\leq l_{2}r\}\cap B_{i}(b_{i},2),B_{i}(b_{i},3))}{r^{nq}}\geq\frac{(2\pi)^{n}}{(C^{\prime})^{n}},\quad 0<r\leq 1,$
as $b\in K\cap\tau_{j_{b}}^{-1}(\mathbb{B}(\mathbf{0},1))$ for some $j_{b}\in\Lambda$. Bi-Lipschitz equivalence of two capacities proved in Lemma 2.9 guarantees a constant $C^{\prime\prime}>0$ satisfying
$\frac{Cap_{\omega}(\{p\in K:d_{\omega}(b,p)\leq R\})}{R^{nq}}\geq C^{\prime\prime},\quad 0<R\leq l_{2},\quad b\in K.$
This is (3.3) in Remark 3.4, which is a uniform density in capacity of $K$ with $nq$.
Now we want to show the Hölder continuity of $V_{K}$. Fix an arbitrary pair $(p,s)$ in $K\times\Lambda$ with $p\in K\cap\tau_{s}^{-1}(\mathbb{B}(\mathbf{0},1))$. Lemma 6.1 guarantees that for
$M_{s}(p):=\|V_{K\cap\bar{B}_{s}(p,2)}^{*}\|_{B_{s}(p,3)}+\|\rho_{2,p,s}\|_{B_{s}(p,3)}+1<\infty,$
$(V_{K\cap\bar{B}_{s}(p,2)}+\rho_{2,p,s})\circ\tau_{s}^{-1}\leq M_{s}(p)u_{\tau_{s}(K\cap\bar{B}_{s}(p,2)),\rho_{2,p,s}\circ\tau_{s}^{-1};\mathbb{B}(\tau_{s}(p),3)}\text{ on }\mathbb{B}(\tau_{s}(p),3).$
The inequality (6.3), comparability (6.2) and inequality (5.7) yield that for each $r\in(0,1]$ and for each $\|z-\tau_{s}(p)\|\leq\delta\leq 1$, as $\tau_{s}(K\cap\bar{B}_{s}(p,r))\subset\bar{\mathbb{B}}(\tau_{s}(p),1)$,
$u_{\tau_{s}(K\cap\bar{B}_{s}(p,r)),\rho_{2,p,s}\circ\tau_{s}^{-1};\mathbb{B}(\tau_{s}(p),3)}(z)$
$\leq(1+\|\rho_{2,p,s}\|_{B_{s}(p,3)})u_{\tau_{s}(K\cap\bar{B}_{s}(p,r));\mathbb{B}(\tau_{s}(p),3)}(z)+\sup_{\tau_{s}(K\cap\bar{B}_{s}(p,r))}\rho_{2,p,s}\circ\tau_{s}^{-1}$
$\leq\frac{1+\|\rho_{2,p,s}\|_{B_{s}(p,3)}}{\inf_{\partial\mathbb{B}(\tau_{s}(p),3)}L_{\tau_{s}(K\cap\bar{B}_{s}(p,r))}}L_{\tau_{s}(K\cap\bar{B}_{s}(p,r))}(z)+\sup_{\tau_{s}(K\cap\bar{B}_{s}(p,r))}\rho_{2,p,s}\circ\tau_{s}^{-1}$
$\leq\frac{1+\|\rho_{2,p,s}\|_{B_{s}(p,3)}}{\inf_{\partial\mathbb{B}(\tau_{s}(p),3)}L_{\bar{\mathbb{B}}(\tau_{s}(p),1)}}\frac{C\delta^{\mu}}{r^{q}}+Dr^{2}$
$=\frac{1+9D}{\log{3}}\frac{C\delta^{\mu}}{r^{q}}+Dr^{2}.$
$\min\{\mu-qt,2t\}$ for $t\in(0,\infty)$ is maximized at $t_{0}:=\frac{\mu}{q+2}\in(0,1)$ as $2t_{0}=\mu-qt_{0}=\frac{2\mu}{q+2}=:\mu^{\prime}<\mu$. Accordingly, for each $\|z-\tau_{s}(p)\|\leq\delta\leq 1$ with $r(\delta):=\delta^{t_{0}}\leq 1$,
$u_{\tau_{s}(K\cap\bar{B}_{s}(p,\delta^{t_{0}})),\rho_{2,p,s}\circ\tau_{s}^{-1};\mathbb{B}(\tau_{s}(p),3)}(z)\leq D^{\prime}\delta^{\mu^{\prime}},\quad D^{\prime}=\frac{(1+9D)C}{\log{3}}+D.$
Then for each $\|z-\tau_{s}(p)\|\leq\delta\leq 1$, since $\rho_{2,p,s}\geq 0$,
$V_{K}\circ\tau_{s}^{-1}(z)$ $\leq(V_{K\cap\bar{B}_{s}(p,2)}+\rho_{2,p,s})\circ\tau_{s}^{-1}(z)$
$\leq M_{s}(p)u_{\tau_{s}(K\cap\bar{B}_{s}(p,2)),\rho_{2,p,s}\circ\tau_{s}^{-1};\mathbb{B}(\tau_{s}(p),3)}(z)$
$\leq M_{s}(p)u_{\tau_{s}(K\cap\bar{B}_{s}(p,\delta^{t_{0}})),\rho_{2,p,s}\circ\tau_{s}^{-1};\mathbb{B}(\tau_{s}(p),3)}(z)\leq M_{s}(p)D^{\prime}\delta^{\mu^{\prime}}.$
For $p\in\tau_{s}^{-1}(\mathbb{B}(\mathbf{0},1))=B_{s}(b_{s},1)$, we have the set inclusion $\bar{B}_{s}(p,2)\supset\bar{B}_{s}(b_{s},1)=\tau_{s}^{-1}(\mathbb{\bar{B}}(\mathbf{0},1))$. Accordingly,
$M_{s}(p)$ $\leq\|V_{K\cap\bar{B}_{s}(b_{s},1)}^{*}\|_{B_{s}(b_{s},4)}+9D+1$
$\leq\sup_{j\in\Lambda:K\cap B_{j}(b_{j},1)\neq\emptyset}\|V_{K\cap\bar{B}_{s}(b_{s},1)}^{*}\|_{B_{s}(b_{s},4)}+9D+1=:D^{\prime\prime}.$
$D^{\prime\prime}$ is a constant independent of the pair $(p,s)\in K\times\Lambda$ of $p\in K\cap\tau_{s}^{-1}(\mathbb{B}(\mathbf{0},1))$. Let $0<\eta\leq 1$ be a Lebesgue number for $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}$ of $(X,d_{\omega})$. Pick $\delta_{0}:=l_{1}\eta/2$. Let $(x,\delta)\in X\times[0,\infty)$ of $d_{\omega}(x,K)\leq\delta\leq\delta_{0}$. There exists $p_{x}\in K$ of $d_{\omega}(x,K)=d_{\omega}(x,p_{x})$. Since $\delta_{0}<\eta$, there exists an index $s_{x}\in\Lambda$ of $\{x,p_{x}\}\subset\tau_{s_{x}}^{-1}(\mathbb{B}(\mathbf{0},1))=B_{s_{x}}(b_{s_{x}},1)$. Then $\|\tau_{s_{x}}(x)-\tau_{s_{x}}(p_{x})\|\leq\delta/l_{1}\leq\delta_{0}/l_{1}\leq 1$,
$V_{K}(x)=V_{K}\circ\tau_{s_{x}}^{-1}(\tau_{s_{x}}(x))\leq M_{s_{x}}(p_{x})D^{\prime}(\delta/l_{1})^{\mu^{\prime}}\leq D^{\prime\prime}D^{\prime}(l_{1})^{-\mu^{\prime}}\delta^{\mu^{\prime}}.$
Therefore, $V_{K}$ is $\mu^{\prime}$-Hölder continuous by Lemma 4.3. ∎
Proof of (iv) $\Rightarrow$ (v) in Theorem 1.2.
Assume that $V_{K}$ is $\nu$-Hölder continuous and $K$ has a uniform density in capacity for constants $0<\nu\leq 1$ and $q_{0}>0$. We will show that $K$ has the local HCP. Fix finite cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of $X$ of holomorphic coordinate balls $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)$ with its refined cover $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$. Let $\varkappa_{0}$ be a constant for this cover for the uniform density in capacity (3.2) of $K$. We use the settings and notations same as those of the proof of (vi) $\Rightarrow$ (iv) in Theorem 1.2: $\bar{B}_{j}(b,r)$, $B_{j}(b,r)$, $b_{j}$, $0<l_{1}\leq 1\leq l_{2}$, $d,D>0$, $\rho_{1,p,j}(\cdot):=d\|\tau_{j}(\cdot)-\tau_{j}(p)\|^{2}$, $\rho_{2,p,j}(\cdot):=D\|\tau_{j}(\cdot)-\tau_{j}(p)\|^{2}$.
Let $C_{0}$ be a bi-Lipschitz constant as in (2.3) such that
$Cap_{\omega}(\cdot\cap B_{j}(b_{j},2))\leq C_{0}Cap(\cdot\cap B_{j}(b_{j},2),B_{j}(b_{j},3)),\quad j\in\Lambda.$
For each $(b,j)\in K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$ and $r\in(0,1]$,
$Cap_{\omega}((K\cap\bar{B}_{j}(b,r))\leq C_{0}\times Cap((K\cap\bar{B}_{j}(b,r))\cap B_{j}(b_{j},2),B_{j}(b_{j},3)).$
Accordingly, for each $(b,j)\in K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
$\frac{Cap((K\cap\bar{B}_{j}(b,r))\cap B_{j}(b_{j},2),B_{j}(b_{j},3))}{r^{q_{0}}}\geq\frac{\varkappa_{0}}{C_{0}},\quad 0<r\leq 1.$
By (3.5), there exists a constant $C_{1}>0$ such that,
(5.8) $\sup_{\bar{\mathbb{B}}(\mathbf{0},3)}L_{\tau_{j}(K\cap\bar{B}_{j}(b,r))}^{*}\leq\frac{C_{1}}{r^{q_{0}}},\quad 0<r\leq 1,\quad j\in\Lambda,\quad b\in K\cap_{j}^{-1}(\mathbb{B}(\mathbf{0},1)).$
By Corollary 5.1-(i), there exist some constants $0<\delta_{1}<2$, $0<A^{\prime}$ such that, for each $(b,j)\in K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
$V_{K\cap\bar{B}_{j}(b,r)}(x)\leq\frac{A^{\prime}}{\varkappa_{0}}\frac{\delta^{\nu}}{r^{q_{0}+2}},\quad d_{\omega}(x,b)\leq\delta\leq\delta_{1},0<r\leq 1.$
The comparability (6.2), inequality (6.3) and Lemma 6.1 in turn imply that, for each $(b,j)\in K\times\Lambda$ with the restrictions
$b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\quad\|z-\tau_{j}(b)\|\leq\delta\leq\delta_{1}/l_{2},\quad 0<r\leq 1,$
with
$m_{j}(b):=\frac{\liminf_{y\to\partial B_{j}(b,2)}\rho_{1,b,j}(y)}{1+\|\rho_{1,b,j}\|_{B_{j}(b,2)}}\equiv\frac{4d}{1+4d}\quad\text{and}$
$c_{j}(b,r):=\sup_{\mathbb{B}(\tau_{j}(b),2)}L_{\tau_{j}(K\cap\bar{B}_{j}(b,r))}\;,$
we have the following chain of inequalities:
$L_{\tau_{j}(K\cap\bar{B}_{j}(b,r))}(z)$ $\leq c_{j}(b,r)u_{\tau_{j}(K\cap\bar{B}_{j}(b,r));\mathbb{B}(\tau_{j}(b),2)}(z)$
$\leq c_{j}(b,r)(u_{\tau_{j}(K\cap\bar{B}_{j}(b,r)),\rho_{1,b,j};\mathbb{B}(\tau_{j}(b),2)}(z)-\inf_{B_{j}(b,2)}\rho_{1,b,j})$
$=c_{j}(b,r)u_{\tau_{j}(K\cap\bar{B}_{j}(b,r)),\rho_{1,b,j};\mathbb{B}(\tau_{j}(b),2)}(z)$
$\leq\frac{c_{j}(b,r)}{m_{j}(b)}(V_{K\cap\bar{B}_{j}(b,r)}(\tau_{j}^{-1}(z))+\rho_{1,b,j}(\tau_{j}^{-1}(z)))$
$\leq\frac{c_{j}(b,r)}{m_{j}(b)}(\frac{A^{\prime}(l_{2}\delta)^{\nu}}{\varkappa_{0}r^{q_{0}+2}}+d\delta^{2})$
$\leq\frac{1+4d}{4d}C_{1}(\frac{A^{\prime}}{\varkappa_{0}}\frac{l_{2}^{\nu}}{r^{2q_{0}+2}}+\frac{d}{r^{q_{0}}})\delta^{\nu}$
$\leq\frac{1+4d}{4d}C_{1}(\frac{A^{\prime}}{\varkappa_{0}}l_{2}^{\nu}+d)\frac{\delta^{\nu}}{r^{2q_{0}+2}}.$
In this chain of inequalities, (5.8) is used to uniformly bound $c_{j}(b,r)$ from above. Therefore, $K$ has local $\nu$-Hölder continuity property of order $2q_{0}+2$. ∎
Remark 5.4.
For a compact $K\subset X$ and $b\in K$, suppose for $\mu_{1}\in(0,1]$, $q_{1}>0$ and some holomorphic coordinate ball $(\Omega_{1},\tau_{1})$ centered at $b\in K$, for some $C^{\prime}_{1}>0$,
$\varpi_{\tau_{1}(K\cap\Omega_{1})\cap\bar{\mathbb{B}}(\mathbf{0},r)}(\mathbf{0},\delta)\leq\frac{C^{\prime}_{1}\delta^{\mu_{1}}}{r^{q_{1}}},\quad 0<\delta\leq 1,\,0<r\leq 1.$
Let $(\Omega_{2},\tau_{2})$ be another holomorphic coordinate ball centered at $b$. There exist $r_{1},r_{2}\in(0,1]$ such that $\tau_{1}(\Omega_{1})\supset\bar{\mathbb{B}}(\mathbf{0},r_{1}),\tau_{2}(\Omega_{2})\supset\bar{\mathbb{B}}(\mathbf{0},r_{2})$. There exist $0<r_{3}\leq r_{1}$ and $l\in(1,\infty)$ such that $\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r_{2}))\supset\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r_{3}))$ and $lr_{3}\leq r_{2}$ and
${\mathbb{B}}(\mathbf{0},lr)\supset\tau_{2}\circ\tau_{1}^{-1}({\mathbb{B}}(\mathbf{0},r))\supset{\mathbb{B}}(\mathbf{0},r/l),\quad r\in(0,r_{3}].$
$\sup_{\bar{\mathbb{B}}(\mathbf{0},r_{3})}L_{\tau_{1}(K\cap\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)))}^{*}\leq\frac{C_{1}^{\prime}r_{3}^{\mu_{1}}}{r^{q_{1}}},\quad 0<r\leq\frac{r_{3}}{l^{3}}.$
From this, by the inequality (3.6),
$\frac{Cap(\tau_{1}(K\cap\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r))),\mathbb{B}(\mathbf{0},r_{3}))}{r^{nq_{1}}}\geq\frac{(2\pi)^{n}}{(C^{\prime}_{1}r_{3}^{\mu_{1}})^{n}},\quad 0<r\leq\frac{r_{3}}{l^{3}}.$
By the monotonicity of the capacity, since $\tau_{2}\circ\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r_{3}))\supset\bar{\mathbb{B}}(\mathbf{0},r_{3}/l)$, we get
$\frac{Cap(\tau_{2}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},lr))),\mathbb{B}(\mathbf{0},r_{3}/l))}{r^{nq_{1}}}\geq\frac{(2\pi)^{n}}{(C^{\prime}_{1}r_{3}^{\mu_{1}})^{n}},\quad 0<r\leq\frac{r_{3}}{l^{3}}.$
By the inequality (3.5), since $l_{3}r_{3}/l^{3}=r_{3}/l^{2}<2r_{3}/(l^{2}+l)<r_{3}/l$, for some constant $c^{\prime}(r_{3},l):=G(2r_{3}/(l^{2}+l))>0$, for $c:=(C^{\prime}_{1}r_{3}^{\mu_{1}})^{n}c^{\prime}(r_{3},l)/(2\pi)^{n}>0$,
$\sup_{\bar{\mathbb{B}}(\mathbf{0},r_{3}/l)}L_{\tau_{2}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},lr)))}^{*}\leq\frac{(C_{1}^{\prime}r_{3}^{\mu_{1}})^{n}c^{\prime}(r_{3},l)}{(2\pi)^{n}r^{nq_{1}}}=\frac{c}{r^{nq_{1}}},\quad 0<r\leq\frac{r_{3}}{l^{3}}.$
By the definition of relative extremal functions in Section 6, for $0<r\leq r_{3}/l^{3}$,
$u_{\tau_{2}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},lr)));\mathbb{B}(\mathbf{0},r_{3}/l)}\circ(\tau_{2}\circ\tau_{1}^{-1})=u_{\tau_{1}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},lr)));\tau_{1}\circ\tau_{2}^{-1}(\mathbb{B}(\mathbf{0},r_{3}/l))}.$
By monotonicity with respect to set inclusion, for $0<r\leq r_{3}/l^{3}$,
$u_{\tau_{1}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},lr)));\tau_{1}\circ\tau_{2}^{-1}(\mathbb{B}(\mathbf{0},r_{3}/l))}\leq u_{\tau_{1}(K\cap\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)));\mathbb{B}(\mathbf{0},r_{3}/l^{2})}.$
By the comparability inequality (6.2), on $\mathbb{B}(\mathbf{0},r_{3}/l^{2})$ and for $0<r\leq r_{3}/l^{3}$,
$u_{\tau_{1}(K\cap\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)));\mathbb{B}(\mathbf{0},r_{3}/l^{2})}\leq(\inf_{\partial\mathbb{B}(\mathbf{0},r_{3}/l^{2})}L_{\tau_{1}(K\cap\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)))})^{-1}L_{\tau_{1}(K\cap\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)))}.$
We can estimate the coefficient on the right-hand side for $0<r\leq r_{3}/l^{3}$ as
$(\inf_{\partial\mathbb{B}(\mathbf{0},r_{3}/l^{2})}L_{\tau_{1}(K\cap\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)))})^{-1}\leq(\inf_{\partial\mathbb{B}(\mathbf{0},r_{3}/l^{2})}L_{\bar{\mathbb{B}}(\mathbf{0},r)})^{-1}=(\log{\frac{r_{3}}{l^{2}r}})^{-1}\leq(\log{l})^{-1}.$
Accordingly, on $\mathbb{B}(\mathbf{0},r_{3}/l^{2})$ and for $0<r\leq r_{3}/l^{3}$,
$u_{\tau_{2}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},lr)));\mathbb{B}(\mathbf{0},r_{3}/l)}\circ(\tau_{2}\circ\tau_{1}^{-1})\leq(\log{l})^{-1}L_{\tau_{1}(K\cap\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)))}.$
Therefore, by the comparability inequality (6.2), for each $\|z\|\leq\delta<r_{3}/l^{2}$ and $0<r\leq r_{3}/l^{3}$, the value $L_{\tau_{2}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},lr)))}\circ(\tau_{2}\circ\tau_{1}^{-1})(z)$ is less than equal to
$(\sup_{\mathbb{B}(\mathbf{0},r_{3}/l)}L_{\tau_{2}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},lr)))}^{*})u_{\tau_{2}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},lr)));\mathbb{B}(\mathbf{0},r_{3}/l)}\circ(\tau_{2}\circ\tau_{1}^{-1})(z),$
which is less than equal to
$\frac{c}{r^{nq_{1}}}\frac{1}{\log{l}}L_{\tau_{1}(K\cap\tau_{1}^{-1}(\bar{\mathbb{B}}(\mathbf{0},r)))}(z)\leq\frac{c}{r^{nq_{1}}}\frac{1}{\log{l}}\frac{C^{\prime}_{1}\delta^{\mu_{1}}}{r^{q_{1}}}.$
Using the variable $t=lr$, we get
$L_{\tau_{2}(K\cap\tau_{2}^{-1}(\bar{\mathbb{B}}(\mathbf{0},t)))}((\tau_{2}\circ\tau_{1}^{-1})(z))\leq\frac{cC^{\prime}_{1}l^{(n+1)q_{1}}}{\log{l}}\frac{\delta^{\mu_{1}}}{r^{(n+1)q_{1}}},\;\|z\|\leq\delta<\frac{r_{3}}{l^{2}},\;t\in(0,\frac{r_{3}}{l^{2}}].$
Therefore, if a compact set $K\subset X$ has the weak local $\mu$-HCP of order $q$ at $b\in K$, then $K$ has local $\mu$-HCP of order $(n+1)q$ at $b\in K$. By the same reason, if $K$ has the weak local $\mu$-HCP of order $q$, then $K$ has local $\mu$-HCP of order $(n+1)q$.
5.3. Proof of the third Theorem
Using Theorem 1.1 and Theorem 1.2, we can prove that a compact subset in $X$ with the weak Hölder continuity property is $(C^{\alpha},C^{\alpha^{\prime}})$-regular (defined in Definition 1.5) for some constants $\alpha$ and $\alpha^{\prime}$ in $(0,1]$ depending on the subset.
Proof of Theorem 1.6.
There exist constants $\mu\in(0,1]$ and $q>0$ such that $K$ has weak local $\mu$-Hölder continuity property of order $q$. Then by Theorem 1.2, $K$ has a uniform density in capacity of order $nq$ and for $\mu^{\prime}:=2\mu/(q+2)$, $V_{K}$ is $\mu^{\prime}$-Hölder continuous.
Mcshane extension [Mc34] extend a Hölder continuous function from a compact subset to $X$ the same Hölder exponent and the same Hölder constant. Since $X$ is compact, each connected component of $X$ has a finite diameter with respect to the Riemannian distance $d_{\omega}$. Therefore, for a given bounded subset $\{\phi_{\gamma}\}_{\gamma\in\Gamma}$ in $C^{\mu^{\prime}}(K)$, each $\phi_{\gamma}$ can be extended to a function $\tilde{\phi}_{\gamma}$ in $C^{\mu^{\prime}}(X)$ so that $\{\tilde{\phi}_{\gamma}\}_{\gamma\in\Gamma}$ is bounded $C^{\alpha}(X)$. Let $\mu^{\prime\prime}:=(\mu^{\prime})^{2}/(\mu^{\prime}+2+q)$. The direct computation gives
$\mu^{\prime\prime}=\frac{4\mu^{2}}{(q+2)(2\mu+(q+2)^{2})}.$
Then by Remark 5.2, for each $\gamma\in\Gamma$, $V_{K,\tilde{\phi}_{\gamma}}$ is $\mu^{\prime\prime}$-Hölder continuous.
Remark 5.2 is a consequence of the proof of Theorem 1.2-(ii). By the proof Theorem 1.2-(ii), we know that $\{V_{K,\tilde{\phi}_{\gamma}}\}_{\gamma\in\Gamma}$ is bounded in $C^{\alpha}(X)$. The reason is as follows. Proposition 4.4 is used in the proof Theorem 1.2-(ii) to extend the Hölder continuity of $V_{K,\tilde{\phi}_{\gamma}}$ on $K$ to the global Hölder continuity. Proposition 4.4 uses Lemma 4.2. In Lemma 4.2, the constant $c_{0}$ depends only on $(X,\omega)$ and the constants $c_{1}$ and $\delta_{1}$ depends only on $(X,\omega,\|V_{K,\tilde{\phi}_{\gamma}}\|_{X})$. By Lemma 2.11, $\|V_{K,\tilde{\phi}_{\gamma}}\|_{X}\leq\|V_{K}\|_{X}+\|\phi_{\gamma}\|_{K}$. Therefore, $\{V_{K,\tilde{\phi}_{\gamma}}\}_{\gamma\in\Gamma}$ is bounded in $C^{\alpha}(X)$. Since $V_{K,\tilde{\phi}_{\gamma}}=V_{K,{\phi}_{\gamma}}$ for each $\gamma\in\Gamma$, we know that $\{V_{K,{\phi}_{\gamma}}\}_{\gamma\in\Gamma}$ is bounded in $C^{\alpha}(X)$. Therefore, $K$ is $(C^{\mu^{\prime}},C^{\mu^{\prime\prime}})$-regular. ∎
6. Appendix
In this section, we give the relation between extremal functions on $X$ and locally defined weighted relative extremal functions in $\mathbb{C}^{n}$. Nguyen in [Ng24] proved this relation for a compact Kähler manifold, and we adjust his proof for a compact Hermitian manifold.
Let $a\in X$. There is a holomorphic coordinate ball $(\Omega,\tau)$ in $X$ centered at $a$, which is the relatively compact restricted chart in another holomorphic chart (if that bigger chart is $(U,f)$, $U$ is open in $X$, $f$ is biholomorphic from $U$ onto an open set in $\mathbb{C}^{n}$, $\Omega\Subset U,f|_{\Omega}=\tau$): $\Omega$ is open in $X$, $\tau$ is biholomorphic with
(6.1) $\tau:\Omega\to\mathbb{B}:=\mathbb{B}(\mathbf{0},1)\subset\mathbb{C}^{n},\quad\tau(\Omega)=\mathbb{B},\quad\tau(a)=\mathbf{0}.$
There exist non-negative bounded smooth strictly psh functions $\rho_{1},\rho_{2}\in PSH(\Omega)\cap C^{\infty}(\Omega)\cap L^{\infty}(\Omega)$ such that $\liminf_{x\to\partial\Omega}\rho_{1}(x)>0$, $\rho_{1}(a)=\rho_{2}(a)=0$ and $dd^{c}\rho_{1}\leq\omega\leq dd^{c}\rho_{2}$. ($\rho_{j}(x):=A_{j}|\tau(x)|^{2},\;x\in\Omega$, for some $A_{j}\in(0,\infty)$ are such functions, as $f|_{\Omega}=\tau$, $\Omega\Subset U$.)
For $E\Subset\mathbb{B}$, the (zero-one) relative extremal function of $E$ on $\mathbb{B}$ is
$u_{E}(z)=u_{E;\mathbb{B}}(z):=\sup\left\{v(z):v\in PSH(\mathbb{B}),v|_{E}\leq 0,v\leq 1\right\},\quad z\in\mathbb{B}.$
[Kl91, Proposition 5.3.3] gives a relation of $L_{F}$ and $u_{F}$ for a compact non-pluripolar set $F$ in $\mathbb{B}$ with $C_{1}:=\inf_{\partial\mathbb{B}}L_{F}\in(0,\infty)$ and $C_{2}:=\sup_{\mathbb{B}}L_{F}^{*}=\sup_{\mathbb{B}}L_{F}\in(0,\infty)$ as
(6.2) $C_{1}\,u_{F}(z)\leq L_{F}(z)\leq C_{2}u_{F}(z),\quad{z\in\mathbb{B}}.$
For a bounded function $\phi:\mathbb{B}\to\mathbb{R}$, the weighted relative extremal function of $(E,\phi)$ on $\mathbb{B}$ is
$u_{E,\phi}(z)=u_{E,\phi;\mathbb{B}}(z):=\sup\{v(z):v\in PSH(\mathbb{B}),\,v|_{E}\leq\phi|_{E},v\leq\phi+1\},\quad z\in\mathbb{B}.$
Like Lemma 2.11-(i), by the definitions, ($\|\cdot\|_{\mathbb{B}}$ denotes the supremum norm on $\mathbb{B}$)
(6.3) $u_{E}+\inf_{\mathbb{B}}\phi\leq u_{E,\phi}\leq(1+\|\phi\|_{\mathbb{B}})u_{E}+\sup_{E}\phi\quad\text{when }\phi\geq 0.$
Like (2.6), for $E_{1}\subset E_{2}\Subset\mathbb{B}$ and $\infty<-\|\phi_{1}\|_{\mathbb{B}}\leq\phi_{1}\leq\phi_{2}\leq\|\phi_{2}\|_{\mathbb{B}}<\infty$,
(6.4) $u_{E_{2},\phi_{1}}\leq u_{E_{1},\phi_{1}}\leq u_{E_{1},\phi_{2}}.$
The following lemma tells a relation between an extremal function on $X$ and the pullback of a weighted relative extremal functions on $\mathbb{B}$ by a chart on $X$.
Lemma 6.1.
Let $\emptyset\neq K$ be a compact non-pluripolar subset of $X$ and $a\in K$. Let $(\Omega,\tau)$ be the holomorphic chart centered at $a$ given in (6.1) and $\rho_{1},\rho_{2}$ be the functions below (6.1). If $K\subset{\Omega}$, then there exist constants $0<M=M(K,\Omega,\|\rho_{2}\|_{\Omega})$ and $0<m=m(\liminf_{x\to\partial\Omega}\rho_{1}(x),\|\rho_{1}\|_{\Omega})$ such that, for $\hat{\rho_{j}}:=\rho_{j}\circ\tau^{-1}$,
$V_{K}\circ\tau^{-1}(z)+\hat{\rho_{2}}(z)\leq M\,u_{\tau(K),\hat{\rho_{2}}}(z),\quad z\in\mathbb{B},$
$m\,u_{\tau(K),\hat{\rho_{1}}}(z)\leq V_{K}\circ\tau^{-1}(z)+\hat{\rho_{1}}(z),\quad z\in\mathbb{B}.$
In fact, the second inequality holds even when $K$ is pluripolar.
Proof.
We use the proof of (6.2) in [Kl91, Proposition 5.3.3]. $\sup_{\Omega}V_{K}^{*}=\|V_{K}^{*}\|_{\Omega}<\infty$ by Proposition 2.4 as $K$ is not pluripolar. Let $v\in PSH(X,\omega)$, $v|_{K}\leq 0$. Then $v+\rho_{2}$ is in $PSH({\Omega})$. Take $M:=\|V_{K}^{*}\|_{\Omega}+\|\rho_{2}\|_{\Omega}+1$. Since $M\geq 1$ and $\rho_{2}\geq 0$,
$(v+\rho_{2})|_{K}\leq{\rho_{2}}|_{K}\leq M\rho_{2}|_{K},\quad v+\rho_{2}\leq M\leq M(1+\rho_{2}).$
These mean $v+\rho_{2}\leq Mu_{\tau(K),\hat{\rho_{2}}}\circ\tau$. Since $v$ was an arbitrary competitor for $V_{K}$, we get the first inequality.
To obtain the second inequality, Let $m^{\prime}:=\liminf_{x\to\partial\Omega}\rho_{1}(x)>0$. Take an open neighborhood $O$ of $K$. Since $O$ is not pluripolar, $V_{O}=V_{O}^{*}\in PSH(X,\omega)$ by Propositions 2.4, 2.5. Let $v\in PSH(\mathbb{B})$ be a competitor for $u_{\tau(K),\hat{\rho_{1}}}$. In other words, $v|_{\tau(K)}\leq\hat{\rho_{1}}|_{\tau(K)}$, $v\leq\hat{\rho_{1}}+1$. Define
$v_{O}|_{\Omega}:=\max\{\frac{m^{\prime}}{1+\|\rho_{1}\|_{\Omega}}v\circ\tau-\rho_{1},V_{O}\}|_{\Omega},\quad v_{O}|_{X\setminus\Omega}:=V_{O}|_{X\setminus\Omega}.$
By Proposition 2.1-(i), as
$\limsup_{x\to\partial\Omega}[\frac{m^{\prime}}{1+\|\rho_{1}\|_{\Omega}}v\circ\tau(x)-\rho_{1}(x)]\leq m^{\prime}-\liminf_{x\to\partial\Omega}\rho_{1}(x)=0\leq V_{O},$
we have $v_{O}\in PSH(X,\omega)$. Let $m:=m^{\prime}/(1+\|\rho_{1}\|_{\Omega})$. Since $m\in(0,1)$, $v\circ\tau|_{K}\leq\rho_{1}|_{K}$ and $\rho_{1}\geq 0$, we have $(mv\circ\tau-\rho_{1})|_{K}\leq 0$. Also, $V_{O}=0$ on $K$. Therefore, $v_{O}\leq V_{K}$, $mv\circ\tau-\rho_{1}\leq V_{K}|_{\Omega}$. Since $v$ was an arbitrary competitor for $u_{\tau(K),\hat{\rho_{1}}}$, the second inequality holds. ∎